### v1

# A Survey on Mixture of Experts

Abstract.
Large language models (LLMs) have garnered unprecedented advancements across diverse fields, ranging from natural language processing to computer vision and beyond. The prowess of LLMs is underpinned by their substantial model size, extensive and diverse datasets, and the vast computational power harnessed during training, all of which contribute to the emergent abilities of LLMs (e.g., in-context learning) that are not present in small models. Within this context, the mixture of experts (MoE) has emerged as an effective method for substantially scaling up model capacity with minimal computation overhead, gaining significant attention from academia and industry. Despite its growing prevalence, there lacks a systematic and comprehensive review of the literature on MoE. This survey seeks to bridge that gap, serving as an essential resource for researchers delving into the intricacies of MoE. We first briefly introduce the structure of the MoE layer, followed by proposing a new taxonomy of MoE. Next, we overview the core designs for various MoE models including both algorithmic and systemic aspects, alongside collections of available open-source implementations, hyperparameter configurations and empirical evaluations. Furthermore, we delineate the multifaceted applications of MoE in practice, and outline some potential directions for future research. To facilitate ongoing updates and the sharing of cutting-edge developments in MoE research, we have established a resource repository accessible at https://github.com/withinmiaov/A-Survey-on-Mixture-of-Experts.
1. Introduction
In the current landscape of artificial general intelligence (AGI), the transformative impact of transformer-based large language models (LLMs) has permeated diverse fields such as natural language processing (Vaswani et al., 2017; Brown et al., 2020; Wei et al., 2022b; Ouyang et al., 2022; Chowdhery et al., 2023; Achiam et al., 2023; Jiang et al., 2024c), computer vision (Riquelme et al., 2021; Liu et al., 2021), and multimodality (Lu et al., 2019; Zhang et al., 2021; Zhou et al., 2022b; Zhu et al., 2023a). Building upon the foundational transformer architecture, LLMs demonstrate extraordinary capabilities, which are attributed to their sheer size, the breadth of data they are trained on, and the significant computational resources invested in their development (Kaplan et al., 2020; Wei et al., 2022a; Yoo et al., 2024). Recognizing a scaling law (Kaplan et al., 2020; Hoffmann et al., 2022) that underpins their evolution, it is imperative to identify and implement efficient methodologies for the sustainable scaling of LLMs.
The concept of mixture of experts (MoE), initially introduced in (Jacobs et al., 1991; Jordan and Jacobs, 1994), has undergone extensive exploration and advancement as evidenced by subsequent studies (Collobert et al., 2001; Rasmussen and Ghahramani, 2001; Shahbaba and Neal, 2009; Eigen et al., 2013; Theis and Bethge, 2015; Deisenroth and Ng, 2015; Aljundi et al., 2017). The emergence of sparsely-gated MoE (Shazeer et al., 2017), particularly within the integration of transformer-based large language models (Lepikhin et al., 2020), has brought new vitality to this three-decade-old technology. The MoE framework is based on a simple yet powerful idea: different parts of a model, known as experts, specialize in different tasks or aspects of the data. With this paradigm, only pertinent experts are engaged for a given input, keeping the computational cost in check while still benefiting from a large pool of specialized knowledge. This scalable and flexible innovation has offered an effective approach for adhering to the scaling law, allowing for increased model capacity without a corresponding surge in computational demands. As depicted in Figure 1, MoE has maintained a robust trajectory of growth, particularly notable in 2024 with the advent of Mixtral-8x7B (Jiang et al., 2024a) and a variety of subsequent industrial-scale LLMs such as Grok-1 (xAI, 2024), DBRX (Databricks, 2024), Arctic (Team, 2024c), DeepSeek-V2 (DeepSeek-AI, 2024), etc.
Despite the increasing popularity and application of MoE models in various domains, the literature has yet to see a survey that thoroughly examines and categorizes the advancements in this area. The most recent review of MoE we could find was presented in September 2022 (Fedus et al., 2022a), predating the pivotal “ChatGPT moment”, which omits the significant advancements that have recently emerged alongside the escalating academic and industrial interest in this domain. This gap in the literature not only hinders the progress of MoE research but also limits the dissemination of knowledge on this topic to a broader audience. Our survey aims to address this deficit by providing a clear and comprehensive overview of MoE with a novel taxonomy that segments recent progress into algorithm, system and application.
Under this taxonomy, we first delve into MoE algorithmic advancements, particularly the prevalent substitution of feed-forward network (FFN) layers with MoE layers in transformer-based LLMs (Lepikhin et al., 2020; Du et al., 2022; Fedus et al., 2022b; Zoph et al., 2022; Jiang et al., 2024a; Xue et al., 2024; DeepSeek-AI, 2024). As each MoE layer integrates multiple FFNs—each designated as an expert—and employs a gating function to activate a select subset of these experts, we explore the design choices of gating function and expert network, alongside collections of available open-source implementations, hyperparameter configurations and empirical evaluations. Furthermore, to underscore the flexibility and versatility of MoE, we extend our analysis beyond the standard integration of MoE into model backbone, and discuss an array of novel MoE-related designs, such as soft MoE with token or expert merging (Puigcerver et al., 2023; Muqeeth et al., 2023; Zhong et al., 2024; Zadouri et al., 2023; Wu et al., 2023a), mixture of parameter-efficient experts (MoPEs) (Wang et al., 2022; Zadouri et al., 2023; Dou et al., 2023b; Gou et al., 2023; Luo et al., 2024; Wu et al., 2024a), training and inference schemes with model transition between dense and sparse (Komatsuzaki et al., 2022; Zhang et al., 2022a; Team, 2023; Xue et al., 2022a; Chen et al., 2022b; Sukhbaatar et al., 2024), and various derivatives (Chen et al., 2023c; Antoniak et al., 2023; Raposo et al., 2024; Xue et al., 2022b; Tan et al., 2023; Choi et al., 2023).
With the gradual convergence of model architecture design in industrial products, system design has emerged as a pivotal factor in enhancing the quality of LLM services. Given the close association of MoE models with machine learning system design, we provide a comprehensive overview of MoE system design, including computation, communication and storage enhancements tailored to address the unique challenges posed by the sparse and dynamic nature of its computational workload. Additionally, we overview the applications of MoE across various domains, including natural language processing, computer vision, recommender system, and multimodal contexts.
The remainder of this survey is organized as follows. Section 2 provides a foundational understanding of MoE, contrasting sparse and dense activation of experts. Section 3 introduces our proposed taxonomy for categorizing MoE advancements. Sections 4, 5, and 6 delve into the algorithmic designs, computing system support, and various applications of MoE models, following the structure outlined in our taxonomy in Figure 3. Finally, in Section 7, we highlight the critical challenges and opportunities for bridging the research-practicality gap, culminating in Section 8 with our conclusions.
2. Background on Mixture of Experts
In transformer-based large language models (LLMs), each mixture of experts (MoE) layer typically consists of a set of “expert networks” , alongside a “gating network” . The role of the gating network, which often takes the form of a linear network with a softmax activation function, is to direct the input to the appropriate expert networks (Shazeer et al., 2017; Fedus et al., 2022b). The MoE layer is strategically placed to select the feed-forward network (FFN) within each Transformer block, typically following the self-attention (SA) sub-layer. This positioning is crucial because the FFN become increasingly computationally demanding as the model scales up. For instance, in the PaLM (Chowdhery et al., 2023) model with the parameter number of 540B, the 90% of these parameters are within its FFN layers.
Formally, each expert network , usually a linear-ReLU-linear network, is parameterized by , accepting the same input and generating an output . In parallel, the gating network , parameterized by and typically consisting of a linear-ReLU-linear-softmax network, yields the output . Based on the design of gating function, MoE layers can be broadly classified into two categories as follows.
2.1. Dense MoE
The dense mixture of experts layer activates all expert networks during each iteration. This strategy has been extensively employed in a range of early works, including those by (Jacobs et al., 1991; Jordan and Jacobs, 1994; Collobert et al., 2001; Rasmussen and Ghahramani, 2001; Shahbaba and Neal, 2009; Eigen et al., 2013; Theis and Bethge, 2015; Deisenroth and Ng, 2015; Aljundi et al., 2017; Ma et al., 2018). Most recent, the dense MoE concept has been revisited by studies such as EvoMoE(Nie et al., 2021), MoLE (Wu et al., 2023b), LoRAMoE (Dou et al., 2023b), and DS-MoE(Pan et al., 2024). The structure of the dense MoE layer is depicted in Figure 2(a). Consequently, the output of the dense MoE layer can be formulated as
where represents the gating value prior to the softmax operation.
2.2. Sparse MoE
While a dense mixture of experts typically yields higher prediction accuracy (Riquelme et al., 2021), it also incurs a significant increase in computational overhead. To address this, Shazeer et al. (Shazeer et al., 2017) introduced the sparsely-gated MoE layer, which is designed to activate only a selected subset of experts during each forward pass. This strategy achieves sparsity by computing a weighted sum of the outputs from only the top- experts, rather than aggregating the outputs from all experts. The structure of the sparse MoE layer is illustrated in Figure 2(b). Building on the framework established by (Shazeer et al., 2017), Equation (2.2) can be modified to reflect the sparsely-gated mechanism as follows:
To explain, function retains only the top- entries of a vector at their original values, while setting all other entries to . Following the operation, those entries assigned become approximately zero. The hyper-parameter is selected based on the specific application, with common choices being (Fedus et al., 2022b; Clark et al., 2022) or (Lepikhin et al., 2020; Du et al., 2022; Rajbhandari et al., 2022; Zoph et al., 2022; Jiang et al., 2024a; Tianwen Wei, 2024). The addition of a noise term is a prevalent strategy for the training of a sparsely-gated MoE layer, fostering exploration among experts and enhancing the stability of MoE training (Shazeer et al., 2017; Fedus et al., 2022b).
Although the sparse gate substantially expands the model’s parameter space without a corresponding increase in computational cost, it can lead to a load balancing issue. Such an issue refers to the uneven distribution of workload across experts, with some being frequently utilized and others seldom or never engaged. To address this, each MoE layer incorporates an auxiliary loss function that promotes an even distribution of tokens across experts within each batch, as described in many studies (Lepikhin et al., 2020; Fedus et al., 2022b; Du et al., 2022; Jiang et al., 2024a; Lieber et al., 2024; Dai et al., 2024; Tianwen Wei, 2024). To formulate this concept, consider a batch of queries , comprising tokens, and experts indexed from to . Following (Lepikhin et al., 2020; Fedus et al., 2022b), the auxiliary load balancing loss for the batch is defined as
where represents the proportion of tokens distributed to expert , while denotes the proportion of the gating probability assigned to expert . To ensure an even distribution of the batch of tokens across the experts, the load-balancing loss function should be minimized. The optimal condition, i.e., , is achieved when each expert receives an equal number of dispatched tokens , and an equal proportion of the gating probability . The balance is thus maintained across all experts, ensuring that the workload is uniformly distributed at all times. Throughout the subsequent sections, unless explicitly stated otherwise, the term “MoE” will refer to “sparse MoE”.
3. Taxonomy of Mixture of Experts
To effectively scale model parameters without a corresponding increase in computational demand, the mixture of experts (MoE) architecture has emerged as a viable solution. MoE leverages a collection of specialized models and a gating mechanism to dynamically select the appropriate “expert networks” for processing a given input. This enables the model to allocate computational resources on an as-needed basis, a concept known as conditional computation. The incorporation of MoE architectures into large language models (LLMs) is now a prevalent practice, allowing these models to achieve significant parameter scale-ups and consequent enhancements in capabilities (Fedus et al., 2022a; Lepikhin et al., 2020; Fedus et al., 2022b; Jiang et al., 2024a; Tianwen Wei, 2024).
forked edges,
for tree=
grow=east,
reversed=true,anchor=base west,
parent anchor=east,
child anchor=west,
base=left,
font=,
rectangle,
draw=line-color,
rounded corners,align=left,
minimum width=2.5em,
s sep=3pt,
inner xsep=2pt,
inner ysep=1pt,
ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center,
where level=1text width=3.2em,font=,,
where level=2text width=3.85em,font=,
where level=3text width=3.48em,font=,where level=4text width=3.8em,font=,[Mixture of Experts (MoE), ver
[Algorithm
[Gating
Function
[Dense
[DS-MoE(Pan et al., 2024), EvoMoE(Nie et al., 2021), MoLE(Wu et al., 2023b), LoRAMoE(Dou et al., 2023b)
,leaf,text width=20.2em]
[Sparse
[Token-Choice
[Shazeer et al.(Shazeer et al., 2017), GShard(Lepikhin et al., 2020), Switch Transformer(Fedus et al., 2022b)
ST-MoE(Zoph et al., 2022), M6-t(Yang et al., 2021), Mod-Squad(Chen et al., 2023a), StableMoE(Dai et al., 2022)
ModuleFormer(Shen et al., 2023c), OpenMoE(Xue et al., 2024), Skywork-MoE (Tianwen Wei, 2024)
Base Layer(Lewis et al., 2021), DSelect-k(Hazimeh et al., 2021), V-MoE(Riquelme et al., 2021), Z-code M3(Kim et al., 2021)
S-Base(Clark et al., 2022), Sentence-level MoE(Kudugunta et al., 2021), NLLB(Costa-jussà et al., 2022), MMoE(Ma et al., 2018)
Task-level MoE(Ye et al., 2022), X-MoE(Chi et al., 2022), Uni-Perceiver-MoE(Zhu et al., 2022)
Mixtral-8x7B(Jiang et al., 2024a), DeepSeekMoE(Dai et al., 2024), Jamba(Lieber et al., 2024), DBRX(Databricks, 2024),
MoA(Zhang et al., 2022b), JetMoE (Shen et al., 2024), Yuan 2.0-M32 (Wu et al., 2024b), DS-MoE(Pan et al., 2024)
,leaf,text width=15em]
[Non-trainable
Token-Choice
[Hash Layer(Roller et al., 2021), THOR(Zuo et al., 2021), DEMix(Gururangan et al., 2022), Task-MoE(Kudugunta et al., 2021)
M2M-100(Fan et al., 2021), Pangu-(Ren et al., 2023)
,leaf,text width=15em]
[Expert-Choice
[Expert-Choice MoE(Zhou et al., 2022a), Brainformers(Zhou et al., 2023)
,leaf,text width=15em]
[Soft
[Token Merging
[Soft MoE(Puigcerver et al., 2023), HOMOE(Dat et al., 2023)
,leaf,text width=15em]
[Expert Merging
[SMEAR(Muqeeth et al., 2023), Lory(Zhong et al., 2024), MoV(Zadouri et al., 2023), Omni-SMoLA(Wu et al., 2023a)
,leaf,text width=15em]
[Experts
[Network
Types
[FFN
[GShard(Lepikhin et al., 2020), Switch Transformer(Fedus et al., 2022b), ST-MoE(Zoph et al., 2022)
Branch-Train-MiX(Sukhbaatar et al., 2024), DS-MoE(Pan et al., 2024), MoEfication(Zhang et al., 2022a)
,leaf,text width=15em]
[Attention
[MoA(Zhang et al., 2022b), JetMoE(Shen et al., 2024), DS-MoE(Pan et al., 2024), ModuleFormer(Shen et al., 2023c)
,leaf,text width=15em]
[Others
[pMoE(Chowdhury et al., 2023), ADVMOE(Zhang et al., 2023), Chen et al.(Chen et al., 2022a), DeepMoE(Wang et al., 2020)
,leaf,text width=15em]
[Hyper-
parameters
[Count
[GShard(Lepikhin et al., 2020), Meta-MoE(Artetxe et al., 2021), GLaM(Du et al., 2022), Mixtral-8x7B(Jiang et al., 2024a)
ST-MoE(Zoph et al., 2022), Swith Transformer(Fedus et al., 2022b), DeepSpeed-MoE(Rajbhandari et al., 2022)
,leaf,text width=15em]
[Size
[GLaM(Du et al., 2022), DeepSeekMoE(Dai et al., 2024), DeepSeek-V2(DeepSeek-AI, 2024), DBRX(Databricks, 2024)
Qwen1.5-MoE(Team, 2024b), LLAMA-MoE (Team, 2023), Mixtral-8x7B(Jiang et al., 2024a)
,leaf,text width=15em]
[Frequency
[ST-MoE(Zoph et al., 2022), OpenMoE(Xue et al., 2024), V-MoE(Riquelme et al., 2021), MoE-LLaVA(Lin et al., 2024)
DeepSeekMoE(Dai et al., 2024), Brainformers(Zhou et al., 2023), Jamba(Lieber et al., 2024)
,leaf,text width=15em]
[Activ. Func.
[ReLU(Glorot et al., 2011), GeLU(Hendrycks and Gimpel, 2016), GeGLU(Shazeer, 2020), SwiGLU(Shazeer, 2020)
,leaf,text width=20.2em]
[Share Expert
[DeepSpeed-MoE(Rajbhandari et al., 2022), NLLB(Costa-jussà et al., 2022), DeepSeekMoE(Dai et al., 2024), OpenMoE(Xue et al., 2024), ScMoE(Cai et al., 2024)
Qwen1.5-MoE(Team, 2024b), MoCLE(Gou et al., 2023), Snowflake-Arctic(Team, 2024c)
,leaf,text width=20.2em]
[PEFT
[FFN
[LoRAMoE(Dou et al., 2023b),
AdaMix(Wang et al., 2022),
MixDA(Diao et al., 2023),
LLaVA-MoLE(Chen et al., 2024)
MixLoRA(Li et al., 2024b)
,leaf,text width=15em]
[Attention
[MoELoRA(Luo et al., 2024),
MoCLE(Gou et al., 2023),
SiRA(Zhu et al., 2023b)
,leaf,text width=15em]
[Transformer
Block
[MoV(Zadouri et al., 2023),
MoLoRA(Zadouri et al., 2023),
UniPELT(Mao et al., 2022),
Omni-SMoLA(Wu et al., 2023a)
MoLA(Gao et al., 2024),
MOELoRA(Liu et al., 2023)
,leaf,text width=15em]
[Every Layer
[MoLE(Wu et al., 2024a)
,leaf,text width=15em]
[Training
& Inference
Scheme
[Original
[Shazeer et al.(Shazeer et al., 2017), GShard(Lepikhin et al., 2020), Switch Transformer(Fedus et al., 2022b), ST-MoE(Zoph et al., 2022)
,leaf,text width=20.2em]
[Dense2Sparse
[RMoE(Wu et al., 2022), Dua et al.(Dua et al., 2022), Sparse Upcycling(Komatsuzaki et al., 2022), DS-MoE(Pan et al., 2024), EvoMoE(Nie et al., 2021)
MoE-LLaVA(Lin et al., 2024), SMoE-Dropout(Chen et al., 2022d), Skywork-MoE(Tianwen Wei, 2024), MoEfication (Zhang et al., 2022a)
MoEBERT(Zuo et al., 2022), LLaMA-MoE(Team, 2023)
,leaf,text width=20.2em]
[Sparse2Dense
[OneS(Xue et al., 2022a), MoE-Pruning(Chen et al., 2022b), ModuleFormer (Shen et al., 2023c), EWA(Huang et al., 2023)
,leaf,text width=20.2em]
[Expert Models
Merging
[Branch-Train-Merge(Li et al., 2022), Branch-Train-MiX(Sukhbaatar et al., 2024), FoE(Wang et al., 2023)
,leaf,text width=20.2em]
[Derivatives
[Lifelong-MoE(Chen et al., 2023c), MoT(Antoniak et al., 2023), MoD(Raposo et al., 2024), WideNet(Xue et al., 2022b), SUT(Tan et al., 2023), SMoP(Choi et al., 2023)
,leaf,text width=25.2em]
[System
[Computation
[FastMoE(He et al., 2021), DeepSpeed-MoE(Rajbhandari et al., 2022), Tutel(Hwang et al., 2023a), SE-MoE(Shen et al., 2022), FasterMoE(He et al., 2022), DeepSpeed-TED(Singh et al., 2023)
HetuMoE(Nie et al., 2022), FlexMoE(Nie et al., 2023), SmartMoE(Zhai et al., 2023), MegaBlocks(Gale et al., 2023),
ScatterMoE(Tan et al., 2024),
PIT(Zheng et al., 2023)
,leaf,text width=25.2em]
[Communication
[DeepSpeed-MoE(Rajbhandari et al., 2022), HetuMoE(Nie et al., 2022), FasterMoE(He et al., 2022), ExFlow(Yao et al., 2024), Tutel(Hwang et al., 2023a), DeepSpeed-TED(Singh et al., 2023)
TA-MoE(Chen et al., 2022c), SE-MoE(Shen et al., 2022),
MPipeMoE (Zhang et al., 2024), SkyWork-MoE(Tianwen Wei, 2024), Lancet(Jiang et al., 2024b), ScMoE(Cai et al., 2024), Arctic(Team, 2024c)
,leaf,text width=25.2em]
[Storage
[SE-MoE(Shen et al., 2022), Pre-gated MoE(Hwang et al., 2023b), EdgeMoE(Yi et al., 2023), MPipeMoE(Zhang et al., 2024)
,leaf,text width=25.2em]
[Application
[NLP
[Shazeer et al.(Shazeer et al., 2017), GShard(Lepikhin et al., 2020), Swith Transformer(Fedus et al., 2022b), Meta-MoE(Artetxe et al., 2021), GLaM(Du et al., 2022), NLLB(Costa-jussà et al., 2022)
Mixtral-8x7B(Jiang et al., 2024a) DeepSpeed-MoE(Rajbhandari et al., 2022), ST-MoE(Zoph et al., 2022), DeepSeekMoE(Dai et al., 2024)
,leaf,text width=25.2em]
[CV
[V-MoE(Riquelme et al., 2021), Swin-MoE(Hwang et al., 2023a), ADVMOE(Zhang et al., 2023), pMoE(Chowdhury et al., 2023)
,leaf,text width=25.2em]
[RecSys
[MMoE(Ma et al., 2018), PLE(Tang et al., 2020), AdaMCT(Jiang et al., 2023b)
,leaf,text width=25.2em]
[MultiModal
[LIMoE(Mustafa et al., 2022), Shen et al.(Shen et al., 2023b), MoCLE(Gou et al., 2023), LLaVA-MoLE(Chen et al., 2024), MoE-LLaVA(Lin et al., 2024), Uni-MoE(Li et al., 2024a)
MM1(McKinzie et al., 2024)
,leaf,text width=25.2em]
For example, the Mixtral 8x7B (Jiang et al., 2024a), introduced by Mixtral AI, shares its foundational architecture with the earlier Mistral 7B (Jiang et al., 2023a), but with a notable difference: each layer comprises eight feed-forward networks (FFN) (i.e., experts). Despite utilizing only 13 billion active parameters, the Mixtral-8x7B demonstrates superior or equivalent performance to the Llama-2-70B (Touvron et al., 2023) and GPT-3.5 (OpenAI, 2022) across various benchmarks. Similarly, the DeepSeek LLM (Bi et al., 2024), developed by DeepSeek, has been extended with an MoE variant known as DeepSeekMoE (Dai et al., 2024). The DeepSeekMoE 16B, while requiring approximately 40% less computation, attains performance on par with the Llama 2 7B (Touvron et al., 2023). The Qwen team has also contributed to this innovative field by developing the Qwen1.5-MoE (Team, 2024b), a smaller MoE model with only 2.7B active parameters that rivals the performance of leading 7B parameter models such as the Mistral 7B (Jiang et al., 2023a) and the Qwen1.5-7B (Team, 2024a).
To assist researchers in navigating the rapidly evolving landscape of LLMs equipped with MoE architectures, we have developed a taxonomy that categorizes these models from three perspectives: algorithm design, system design, and application. Figure 3 showcases our taxonomy alongside several representative studies. In the following sections, we will provide a comprehensive and in-depth analysis of each category within our taxonomy.
4. Algorithm Design of Mixture of Experts
4.1. Gating Function
The gating function (also known as the routing function or router), which stands as a fundamental component of all the MoE architectures, orchestrates the engagement of expert computations and the combination of their respective outputs. We categorize this mechanism into three distinct types Based on the processing methodologies of each input, we categorize the gating mechanism into three distinct types: sparse, which activates a subset of experts; dense, which activates all experts; and soft, which encompasses fully-differentiable approaches including input token merging and expert merging.
4.1.1. Sparse
The sparse gating functions activate a selected subset of experts for processing each individual input token, which can be considered as a form of conditional computation (Bengio et al., 2013; Davis and Arel, 2013; Almahairi et al., 2016). The gating functions have been studied extensively, which may be trained by various forms of reinforcement learning and back-propagation, making binary or sparse and continuous, stochastic or deterministic gating decisions (Eigen et al., 2013; Bengio et al., 2015; Rosenbaum et al., 2017, 2019; Clark et al., 2022). Shazeer et al. (Shazeer et al., 2017) pioneered a differentiable heuristic with auxiliary load balancing losses, in which the outputs from expert computations are weighted by their selection probabilities. This introduces a differentiable aspect to the gating process, thereby facilitating the derivation of gradients that can guide the gating function’s optimization. This paradigm has subsequently become predominant in the realm of MoE research. Due to its selection of experts for each input token, this method can be recognized as a gating function with token choice.
Token-Choice Gating. Shazeer et al. (Shazeer et al., 2017) posited the necessity of gating inputs to the top- experts, with , to enhance the efficacy of MoE. The rationale behind this approach is that by simultaneously consulting multiple experts for a given input, the network can effectively weigh and integrate their respective contributions, thereby improving performance. To accommodate the scalability to thousands of experts within a MoE layer, they employ a two-level hierarchical MoE to reduce the branching factor in the context of a large expert count. Subsequent research has largely affirmed that increasing the value of k enhances performance, which has led to the widespread adoption of this top- strategy with . Notwithstanding, the Switch Transformer model (Fedus et al., 2022b) has shown that a top-1 gating strategy (as illustrated in Figure 4 (a)) can also yield competitive results, a finding that has been substantiated and adopted by later studies (Clark et al., 2022). Furthermore, M6-t (Yang et al., 2021) proposed a novel variation of the top-1 gating called expert prototyping, which organizes experts into k groups and then applies top-1 gating in each group. Their experimental results show the training and downstream perplexity of a 16-layer model in order of best to worst: expert prototyping with 4 top-1 gating, 1 top-4 gating, 1 top-16 gating, 1 top-1 gating.
Auxiliary Loss for Token-Choice Gating. Token-choice gating algorithms frequently incorporate an auxiliary loss during training to promote equitable token distribution across experts. Table 1 shows prevalent auxiliary loss functions leveraged in the field. Shazeer et al. (Shazeer et al., 2017) quantify the importance of an expert in relation to a training batch via the batchwise sum of the gate values for that expert. They define an additional loss , added to the overall loss function for the model. This loss, which is equal to the square of the coefficient of variation of the set of importance values and multiplied by a hand-tuned scaling factor , encourages all experts to have equal importance. Although promotes balance in importance, it does not guarantee an even distribution of training examples among experts, which can lead to execution inefficiencies on distributed computing environments. To address this, they introduce a second loss to ensure balanced loads. Building on this foundation, GShard (Lepikhin et al., 2020) define a new differentiable auxiliary loss using a differentiable approximation (the dot-product of mean gates and mean gating decisions per expert), as detailed in Section 2.2. Switch Transformers (Fedus et al., 2022b) and many other subsequent studies (Du et al., 2022; Jiang et al., 2024a; Databricks, 2024; Lieber et al., 2024) have embraced this design, and enhancements (Dai et al., 2024; DeepSeek-AI, 2024; Tianwen Wei, 2024) have been made to cater to diverse requirements. Nevertheless, ST-MoE (Zoph et al., 2022) identified limitations with , particularly at larger scales, leading to unreliable training outcomes. To mitigate this, it introduces the integration of router z-loss , improving training stability without quality degradation by penalizing large logits entering the gating network. Since this loss encourages absolute magnitude of values to be smaller, roundoff errors are reduced, which can be quite impactful for exponential functions such as the gating. Additionally, Mod-Squad (Chen et al., 2023a) posits the difficulty of training multi-task models under such an expert-balancing loss, which may inadvertently force experts to set parameters on conflicting tasks or hinder the potential synergies from parameter sharing across complementary tasks. Instead, it proposes to maximize the mutual information (MI) between experts and tasks to build task-expert alignment. Differently, ModuleFormer (Shen et al., 2023c) proposes to maximize the Mutual Information between experts and tokens. Furthermore, DS-MoE (Pan et al., 2024) extends the application of , calibrating different weightings , in Mixture-of-Attention (MoA, as illustrated in Figure 5 (a)) and FFN MoE modules of different size models.
| Reference | Auxiliary Loss | Coefficient |
| Shazeer et al.(Shazeer et al., 2017), V-MoE(Riquelme et al., 2021) | , | |
| GShard(Lepikhin et al., 2020), Switch-T(Fedus et al., 2022b), GLaM(Du et al., 2022), Mixtral-8x7B(Jiang et al., 2024a), DBRX(Databricks, 2024), Jamba(Lieber et al., 2024), DeepSeekMoE(Dai et al., 2024), DeepSeek-V2(DeepSeek-AI, 2024), Skywork-MoE(Tianwen Wei, 2024) | ||
| ST-MoE(Zoph et al., 2022), OpenMoE(Xue et al., 2024), MoA(Zhang et al., 2022b), JetMoE (Shen et al., 2024) | , | |
| Mod-Squad(Chen et al., 2023a), Moduleformer(Shen et al., 2023c), DS-MoE(Pan et al., 2024) |
Expert Capacity for Token-Choice Gating. In conjunction with load balancing via auxiliary loss, GShard (Lepikhin et al., 2020) incorporates an expert capacity limit, defining a threshold for the number of tokens an expert can process. This can lead to token overflow, where excess tokens are not processed by the designated expert. GShard also proposes a random routing mechanism that selects a secondary expert with a probability proportional to its weight, under the intuition that the contribution of a secondary expert can be negligible, given that the output is a weighted average and the secondary weight is typically small. For the task of image classification with Vision Transformer (ViT) models, Riquelme et al. (Riquelme et al., 2021) enhance the top- gating strategy with Batch Prioritized Routing (BPR), which assigns priority based on higher gating scores rather than the sequence order of tokens. Zoph et al. (Zoph et al., 2022) have demonstrated the efficacy of BPR in the context of MoE language models. Kim et al. (Kim et al., 2021) suggest randomizing token prioritization within sequences to mitigate routing bias towards early-positioned tokens. OpenMoE (Xue et al., 2024) provides a comprehensive analysis of gating mechanisms, highlighting the “Drop-towards-the-End” phenomenon whereby tokens later in a sequence are at greater risk of being dropped due to experts reaching their maximum capacity limits, an issue that is exacerbated in instruction-tuning datasets. Moreover, OpenMoE identifies a tendency within MoE systems to route tokens based on token-level semantic similarities, leading to “Context-independent Specialization”. Additionally, this token ID routing specialization is established early in pre-training and remains largely fixed, resulting in a consistent pattern of token processing by the same experts throughout training, a phenomenon referred to as “Early Routing Learning”.
Other Advancements on Token-Choice Gating. Despite the implementation of gating heuristics and auxiliary expert-balancing loss functions aimed at achieving a balanced workload distribution among experts, the issue of load imbalance persists as a prevalent challenge within MoE architectures. To solve it, the Balanced Assignment of Sparse Experts (BASE) layer, as conceptualized by Lewis et al. (Lewis et al., 2021) and illustrated in Figure 4 (b), re-envisions the token-to-expert allocation process by casting it as a linear assignment problem, aiming to maximize the token-expert affinities under the constraints that each expert is assigned an equal quantity of tokens. Subsequently, Clark et al. (Clark et al., 2022) introduce a variant of the BASE layer, termed S-BASE, using an optimal transport formulation. Additionally, they devise a reinforcement learning based gating algorithm employing top-1 routing, with the reward function defined as the negative cross-entropy of the predicted token. In addressing the discrete optimization challenge of gating function that can lead to convergence and statistical performance issues when training with gradient-based methods, Hazimeh et al. (Hazimeh et al., 2021) introduce DSelect-k. which is a smooth version of the top- gating algorithm that improves over standard top- gating. This method constitutes a refined version of the top- gating algorithm, featuring enhanced smoothness properties that yield improvements over the conventional top- gating approach. Kudugunta et al. (Kudugunta et al., 2021) diverge from the prevalent token-level gating strategies by introducing a sentence-level gating mechanism. This approach involves generating a sentence representation by averaging the tokens within a sequence and subsequently routing it to an expert. Chi et al.(Chi et al., 2022) observe that prevailing gating mechanisms tend to push hidden representations clustering around expert centroids, implying a trend toward representation collapse, which in turn harms model performance. To counteract this issue, they project hidden vectors into a lower-dimensional space before gating and implement L2 normalization for both token representations and expert embeddings, thus calculating gating scores within a low-dimensional hypersphere. Skywork-MoE (Tianwen Wei, 2024) proposes two innovative techniques: gating logit normalization, which improves expert diversification, and adaptive auxiliary loss coefficients, which provides layer-specific adjustment of auxiliary loss coefficients. Yuan 2.0-M32 (Wu et al., 2024b) proposes a new router network, Attention Router (as illustrated in Figure 4 (e)), which implements a more efficient selection of experts and yields an enhancement in model accuracy over classical linear router network.
Non-trainable Token-Choice Gating. The dynamic training of gating functions within MoE models is standard practice; however, some research has ventured into the realm of non-trainable token-choice gating mechanisms. The most significant benefit of non-trainable token-choice gating is that no additional gating network parameters are required and the full load balancing can be achieved through specific gating mechanisms. The Hash Layer (Roller et al., 2021) utilizes a random fixed gating approach by hashing the input token, achieving competitive results without the necessity of training the gating network. The load balancing is facilitated by the selection of hash functions prior to training, which can equitably distribute token batches. Zuo et al. (Zuo et al., 2021) introduces THOR, an algorithm that randomly allocates two experts to each input during training and inference with a consistency regularized loss promoting consistent predictions. Gururangan et al. (Gururangan et al., 2022) propose the DEMix model, which explicitly assigns distinct experts to discrete pre-training domains, with domain matching being employed to select experts corresponding to the training inputs. Given the potential suboptimality of domain categorization and its limited scope in encompassing test-time domains, a single domain expert selection could undermine the model’s generalizability. To address this, DEMix adopts a parameter-free probabilistic method that dynamically estimates the domain-weighted mixture at inference. Kudugunta et al. (Kudugunta et al., 2021) explore task-level gating incorporating prior knowledge tags, and similarly, M2M-100 model (Fan et al., 2021) utilizes explicit language-specific sublayers with deterministically routing input tokens based on their language. Building upon the aforementioned non-trainable gating strategies—random gating and domain mapping—PanGu- (Ren et al., 2023) presents the Random Routed Experts (RRE) mechanism. As illustrated in Figure 4 (c), this approach initially routes tokens to a domain-specific expert group, followed by a random selection within that group.
In contrast to explicit language-specific expert selection, NLLB (Costa-jussà et al., 2022) leverages trainable gating to manage multilingual machine translation tasks, outperforming the M2M-100 approach (Fan et al., 2021). Addressing task interference in generalist models, Zhu et al. (Zhu et al., 2022) introduce the Conditional MoE, which augments MoE with trainable gating by integrating conditional information at various levels, such as token-level, context-level, modality-level, task-level, and predefined token attributes. Ye et al. (Ye et al., 2022) further investigate the incorporation of trainable gating at task-level MoE. Additionally, STABLEMOE (Dai et al., 2022) identifies a challenge with existing learning-to-route MoE methods: the phenomenon of gating fluctuation. To counter this, STABLEMOE employs a two-stage training process. The first stage focuses on acquiring a balanced and cohesive gating strategy, which is then distilled into a lightweight gate function, decoupled from the backbone model. Subsequently, the second stage leverages the distilled gate for token-to-expert assignments and freezes it to ensure a stable gating strategy throughout further training.
Expert-Choice Gating. Zhou et al. (Zhou et al., 2022a) propose an inversion of the conventional token-choice gating paradigm, wherein each expert selects the top- tokens they will process, as illustrated in Figure 4 (d). This approach circumvents the necessity for auxiliary load balancing losses during training, ensuring a uniform distribution of tokens across experts. However, this method may result in uneven token coverage, with some tokens potentially being processed by multiple experts or not at all. Despite this, the technique demonstrates strong empirical performance and offers an adaptive computational interpretation where the model can implicitly apply more computation to certain tokens. The effectiveness of expert-choice gating is further validated by Zhou et al. in their subsequent Brainformers study (Zhou et al., 2023). Additionally, Komatsuzaki et al. (Komatsuzaki et al., 2022) integrate the expert-choice gating strategy within the Vision Transformer and adapt it for the encoder in T5 models, while maintaining token-choice gating for the T5 decoder.
4.1.2. Dense
In Section 2.1, we discuss the enduring relevance of dense MoE, which activates all the experts for each input process. This dense paradigm continues to inform current innovations in MoE training and inference methodologies, as elaborated in Section 4.4.1. While sparse activation of experts, as a trade-off, may yield efficiency gains at the expense of some performance when compared to a densely activated MoE with an equivalent number of total parameters (Dai et al., 2024; Pan et al., 2024; Shen et al., 2023c), it represents a strategic adjustment to balance computational demands with model capability. Notably, dense activation performs well in the context of LoRA-MoE fine-tuning, where the computational overhead of LoRA experts is comparatively low. This approach enables the effective and flexible integration of multiple LoRAs across a variety of downstream tasks. It preserves the generative capabilities of the original pre-trained model and maintains the unique characteristics of individual LoRAs for each task (Wu et al., 2023b; Dou et al., 2023b).
4.1.3. Soft
Deciding the allocation of appropriate experts to each input token pose the fundamental discrete optimization challenge for sparse MoE. This often necessitates heuristic auxiliary losses to ensure balanced expert engagement and to minimize unassigned tokens. These issues become more pronounced in scenarios involving out-of-distribution data, such as small inference batches, novel inputs, or during transfer learning. Similar to dense MoE, the soft MoE approach maintains full differentiability by leveraging all the experts for processing each input, thus avoiding issues inherent to discrete expert selection. We distinguish soft MoE from dense MoE to highlight the characteristic that mitigates computational demands through the gating-weighted merging of input tokens or experts.
Token Merging. Puigcerver et al. (Puigcerver et al., 2023) proposed the Soft MoE, which eschews the conventional sparse and discrete gating mechanism in favor of a soft assignment strategy that merges tokens. This method computes several weighted averages of all tokens, with weights depending on both tokens and experts, and processes each aggregate with its respective expert. Their experimental results in image classification demonstrate that soft MoE enhances the stability of gating function training and inherently maintains balance. HOMOE (Dat et al., 2023) follows the design of Soft MoE and combines it with Hopfield network to address the the challenges of Compositional Zero-Shot Learning tasks. Yet, merging input tokens complicates its application in auto-regressive decoders, as future tokens required for averaging are inaccessible during inference.
Expert Merging. In contrast to the merging of input tokens, Muqeeth et al. (Muqeeth et al., 2023) introduced the Soft Merging of Experts with Adaptive Routing (SMEAR) framework, which circumvents discrete gating by merging all the experts’ parameters through a weighted average, as illustrated in Figure 4 (f). They argue that conventional sparse MoE models often fail to match the performance of their parameter-matched dense counterparts or those utilizing non-learned heuristic gating functions, potentially due to flawed gradient estimation methods for training modules with non-differentiable, discrete gating decisions. By processing the input tokens through a single merged expert, SMEAR does not incur a significant increase in computational costs and enables standard gradient-based training. Empirical evaluations on T5-GLUE and ResNet-DomainNet benchmarks reveal that SMEAR-equipped models surpass those with metadata-based (Gururangan et al., 2022; Kudugunta et al., 2021) or gradient-estimated learning gating strategies. On ResNet-DomainNet, SMEAR achieved a 1.5% higher average accuracy than Soft MoE (Puigcerver et al., 2023) with single “slot” per expert, at the expense of a near 10% reduction in throughput. Subsequent contributions by Zhong et al. (Zhong et al., 2024) argue that SMEAR’s demonstrated advantages are confined to downstream fine-tuning on classification tasks. They present Lory, an innovative approach for scaling such expert merging architectures to auto-regressive language model pre-training. Lory (Zhong et al., 2024) introduces a causal segment routing strategy, conducting expert merging at the segment level while maintaining the auto-regressive nature of language models. Furthermore, it employs similarity-based data batching to direct expert specialization in particular domains or topics. Lory’s empirical validation on LLaMA models showcases significant improvements over parameter-matched dense models in terms of perplexity (by 13.9%) and on diverse downstream tasks (by 1.5%-11.1%), highlighting the potential of fully-differentiable MoE architectures for language model pre-training and encouraging further investigation in this area. In addition, expert merging methods have demonstrated efficacy in parameter-efficient fine-tuning (PEFT) MoE contexts. Zadouri et al. (Zadouri et al., 2023) substantiate that soft merging of experts significantly outperforms sparse gating mechanisms (top-1, top-2) in the T5 models (Raffel et al., 2020) fine-tuning with the MoV-10 setting of 10 (IA)3 vector expert. Wu et al. (Wu et al., 2023a) propose Omni-SMoLA, an architecture leveraging the soft method to mix multimodal low-rank experts, improving the generalist performance across a broad range of generative vision-language tasks.
4.2. Experts
In this section, we delineate the architecture of expert networks within MoE framework, following our discussion on the gating function that orchestrates the activation of these experts.
4.2.1. Network Types
Since the initial integration of MoE into transformer architectures (Lepikhin et al., 2020; Fedus et al., 2022b; Zoph et al., 2022), MoE has served as a substitute for Feed-Forward Network (FFN) modules within these models. Typically, each expert within a MoE layer replicates the architecture of the FFN it replaces. This paradigm, wherein FFNs are utilized as experts, remains predominant, and subsequent refinements will be expounded upon in Sections 4.2.2 to 4.2.4.
Feed-Forward Network. As discussed in existing work (Sukhbaatar et al., 2024), the predilection for leveraging MoE in the context of FFNs is rooted in the hypothesis that self-attention layers exhibit lower sparsity and less domain specificity than FFN layers. Pan et al. (Pan et al., 2024) provide empirical support for this, revealing marked sparsity in FFN layers compared to self-attention layers, through their analysis of downstream Wikitext tasks using their pre-trained DS-MoE models. Their results indicate a mere 20% active expert engagement in FFN layers, in contrast to the 80% observed within self-attention layers. In earlier investigation of FFN computational patterns, Zhang et al. (Zhang et al., 2022a) observe that most inputs only activate a small proportion of neurons of FFNs, thus corroborating the inherent sparsity of FFNs.
Attention. While the focus of MoE research has predominantly been on FFN layers within the Transformer architecture, Zhang et al. (Zhang et al., 2022b) introduce the Mixture of Attention Heads (MoA), an innovative architecture that combines multi-head attention layers with MoE to further enhance performance and restrain computational cost. As delineated in Figure 5 (a), MoA employs two sets of experts, one for query projection and one for output projection, which are selected the same indices of experts through a common gating network. To reduce computational complexity, MoA shares the key () and value () projection weights across attention experts, with experts differentiated only by their respective query () and output () projection weights, allowing for shared pre-computation of key () and value () sequences. Subsequent work such as DS-MoE (Pan et al., 2024), JetMoE (Shen et al., 2024), and ModuleFormer (Shen et al., 2023c) follows the design of MoA and further refines the combination of MoE and attention layer.
Others. In addition to the aforementioned expert network types, researchers have explored the use of Convolutional Neural Network (CNN) as expert (Chowdhury et al., 2023; Zhang et al., 2023; Chen et al., 2022a; Wang et al., 2020; Gross et al., 2017). Moreover, recent endeavors that integrate Parameter-Efficient Fine-Tuning (PEFT) techniques with MoE, such as employing Low-Rank Adaptation (LoRA) (Hu et al., 2021) as expert, have shown promising results, which are discussed in Section 4.3.
| Reference | Models | Expert Count (Activ./Total) | #L | #H | Placement Frequency | Activation Function | Share Expert Count | ||||
| GShard (Lepikhin et al., 2020) (2020) | 600B | 2/2048 | 1024 | 8192 | 36 | 16 | 128 | 1/2 | ReLU | 0 | |
| 200B | 2/2048 | 1024 | 8192 | 12 | 16 | 128 | 1/2 | ReLU | 0 | ||
| 150B | 2/512 | 1024 | 8192 | 36 | 16 | 128 | 1/2 | ReLU | 0 | ||
| 37B | 2/128 | 1024 | 8192 | 36 | 16 | 128 | 1/2 | ReLU | 0 | ||
| Switch (Fedus et al., 2022b) (2021) | 7B | 1/128 | 768 | 2048 | 12 | 12 | 64 | 1/2 | GEGLU | 0 | |
| 26B | 1/128 | 1024 | 2816 | 24 | 16 | 64 | 1/2 | GEGLU | 0 | ||
| 395B | 1/64 | 4096 | 10240 | 24 | 64 | 64 | 1/2 | GEGLU | 0 | ||
| 1571B | 1/2048 | 2080 | 6144 | 15 | 32 | 64 | 1 | ReLU | 0 | ||
| GLaM (Du et al., 2022) (2021) | 0.1B/1.9B | 2/64 | 768 | 3072 | 12 | 12 | 64 | 1/2 | GEGLU | 0 | |
| 1.7B/27B | 2/64 | 2048 | 8192 | 24 | 16 | 128 | 1/2 | GEGLU | 0 | ||
| 8B/143B | 2/64 | 4096 | 16384 | 32 | 32 | 128 | 1/2 | GEGLU | 0 | ||
| 64B/1.2T | 2/64 | 8192 | 32768 | 64 | 128 | 128 | 1/2 | GEGLU | 0 | ||
| DeepSpeed-MoE (Rajbhandari et al., 2022) (2022) | 350M/13B | 2/128 | 1024 | 24 | 16 | 64 | 1/2 | GeLU | 0 | ||
| 1.3B/52B | 2/128 | 2048 | 24 | 16 | 128 | 1/2 | GeLU | 0 | |||
| PR-350M/4B | 2/32-2/64 | 1024 | 24 | 16 | 64 | 1/2, 10L-32E, 2L-64E | GeLU | 1 | |||
| PR-1.3B/31B | 2/64-2/128 | 2048 | 24 | 16 | 128 | 1/2, 10L-64E, 2L-128E | GeLU | 1 | |||
| ST-MoE (Zoph et al., 2022) (2022) | 0.8B/4.1B | 2/32 | 1024 | 2816 | 27 | 16 | 64 | 1/4, add extra FFN | GEGLU | 0 | |
| 32B/269B | 2/64 | 5120 | 20480 | 27 | 64 | 128 | 1/4, add extra FFN | GEGLU | 0 | ||
| Mixtral (Jiang et al., 2024a) (2023) | 13B/47B | 2/8 | 4096 | 14336 | 32 | 32 | 128 | 1 | SwiGLU | 0 | |
| 39B/141B | 2/8 | 6144 | 16384 | 56 | 48 | 128 | 1 | SwiGLU | 0 | ||
| LLAMA-MoE (Team, 2023) (2023) | 3.0B/6.7B | 2/16 | 4096 | 11008 | 688 | 32 | 32 | 128 | 1 | SwiGLU | 0 |
| 3.5B/6.7B | 4/16 | 4096 | 11008 | 688 | 32 | 32 | 128 | 1 | SwiGLU | 0 | |
| 3.5B/6.7B | 2/8 | 4096 | 11008 | 1376 | 32 | 32 | 128 | 1 | SwiGLU | 0 | |
| DeepSeekMoE (Dai et al., 2024) (2024) | 0.24B/1.89B | 8/64 | 1280 | - | 9 | 10 | 128 | 1 | SwiGLU | 1 | |
| 2.8B/16.4B | 8/66 | 2048 | 10944 | 1408 | 28 | 16 | 128 | 1, except 1st layer | SwiGLU | 2 | |
| 22B/145B | 16/132 | 4096 | - | 62 | 32 | 128 | 1, except 1st layer | SwiGLU | 4 | ||
| OpenMoE (Xue et al., 2024) (2024) | 339M/650M | 2/16 | 768 | 3072 | 12 | 12 | 64 | 1/4 | SwiGLU | 1 | |
| 2.6B/8.7B | 2/32 | 2048 | 8192 | 24 | 24 | 128 | 1/6 | SwiGLU | 1 | ||
| 6.8B/34B | 2/32 | 3072 | 12288 | 32 | 24 | 128 | 1/4 | SwiGLU | 1 | ||
| Qwen1.5-MoE (Team, 2024b) (2024) | 2.7B/14.3B | 8/64 | 2048 | 5632 | 1408 | 24 | 16 | 128 | 1 | SwiGLU | 4 |
| DBRX (Databricks, 2024) (2024) | 36B/132B | 4/16 | 6144 | 10752 | 40 | 48 | 128 | 1 | SwiGLU | 0 | |
| Jamba (Lieber et al., 2024) (2024) | 12B/52B | 2/16 | 4096 | 14336 | 32 | 32 | 128 | 1/2, 1:7 Attention:Mamba | SwiGLU | 0 | |
| Skywork-MoE (Tianwen Wei, 2024) (2024) | 22B/146B | 2/16 | 4608 | 12288 | 52 | 36 | 128 | 1 | SwiGLU | 0 | |
| Yuan 2.0-M32 (Wu et al., 2024b) (2024) | 3.7B/40B | 2/32 | 2048 | 8192 | 24 | 16 | 256 | 1 | SwiGLU | 0 |
4.2.2. Hyperparameters
The scale of sparse MoE models is governed by several critical hyperparameters that extend beyond those of dense transformer models. These include (1) the count of experts per MoE layer, (2) the size of each expert, and (3) the placement frequency of MoE layers throughout the model. The selection of these hyperparameters is crucial, as it profoundly influences model performance and computational efficiency across various tasks. Optimal hyperparameter choices are thus contingent upon the specific application requirements and the constraints of the computational infrastructure. Our subsequent analysis, informed by the exemplified models listed in Table 2, explores these hyperparameter decisions in depth. Meanwhile, we enumerate some recent open-source models, detailing their number of parameters and benchmark performance in Table 3.
Expert Count. Initial investigations employing thousands of experts per layer yielded impressive gains in pre-training and translation quality (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022b). Nonetheless, the quality of sparse MoE models is disproportionately reduced under domain shift (Artetxe et al., 2021) or when fine-tuning on diverse task distributions (Fedus et al., 2022b). GLaM (Du et al., 2022) adopts a configuration of 64 experts, guided by their findings that a 64-expert setup with top-2 gating strikes an optimal balance between execution efficiency and performance across zero-shot, one-shot, and few-shot scenarios. Reflecting this trend, more recent sparse MoE models (Zoph et al., 2022; Jiang et al., 2024a; Xue et al., 2024; Team, 2024b; Databricks, 2024; Tianwen Wei, 2024; Lieber et al., 2024; Wu et al., 2024b) commonly utilize no more than 64 experts. Additionally, DeepSpeed-MoE (Rajbhandari et al., 2022) adopts a Pyramid-MoE approach, positioning MoE layers with a larger expert count towards the network’s end.
Expert Size. To scale the model effectively, GLaM (Du et al., 2022) prioritizes the expansion of the intermediate hidden dimension per expert while standardizing the expert count at 64, a strategy that often requires the implementation of tensor parallelism across multiple accelerators to maintain computational efficiency (Du et al., 2022; Rajbhandari et al., 2022; Fedus et al., 2022b). From this period forward, MoE models (Zoph et al., 2022; Jiang et al., 2024a; Databricks, 2024; Tianwen Wei, 2024) typically featured larger expert dimensions. Differently, DeepSeekMoE (Dai et al., 2024; DeepSeek-AI, 2024) introduces the concept of fine-grained expert segmentation by subdividing the intermediate hidden dimension of FFN expert, while preserving the overall parameter count. Specifically, DeepSeekMoE-145B employs a reduced intermediate hidden dimension at one-eighth that of its dense FFN counterpart, increasing both the number of experts (from 16 to 128) and the number of active experts (from top-2 to top-16) by a factor of eight. They believe that this strategy not only refines the decomposition of knowledge across experts, facilitating more precise learning, but also enhances the flexibility of expert activation combinations, allowing for more specialized and targeted knowledge capture. Qwen1.5-MoE (Team, 2024b) and DBRX (Databricks, 2024) adopt a similar fine-grained expert segmentation strategy. From the results of LLAMA-MoE, which evenly splits the parameters of FFNs into non-overlapping experts to construct the MoE models with same parameter count, activating 4 of 16 expert performs slightly better than activating 2 of 8 expert. Results from LLAMA-MoE (Team, 2023), which allocates dense FFN parameters across non-overlapping experts to maintain a consistent parameter count, indicate that activating 4 out of 16 experts with a dimensionality of marginally outperforms the activation of 2 out of 8 experts with .
Frequency of MoE Layers. Sparse MoE models typically evolve from dense architectures by interspersing MoE layers in place of the dense FFN layers at regular intervals. Although a higher frequency of MoE layers can enlarge the model size, it also introduces greater system overhead. In practice, most MoE models features alternate FFN replacement (1/2) with MoE layers (Lepikhin et al., 2020; Du et al., 2022; Artetxe et al., 2021; Rajbhandari et al., 2022). Nevertheless, variations exist, with some models incorporating MoE layers every fourth layer (1/4) (Zoph et al., 2022; Xue et al., 2024) or in every layer (1/1) (Fedus et al., 2022b; Dai et al., 2024). Following the introduction of Mixtral 8x7B (Jiang et al., 2024a), the trend seems to shift towards placing MoE in every layer of the model, with a common choice of only 8 or 16 experts mirroring the dimensionality of a dense FFN (Dai et al., 2024; Team, 2024b; Databricks, 2024; Tianwen Wei, 2024).
Research into the optimal configuration of MoE layers has been extensive. V-MoE (Riquelme et al., 2021) employs MoE in the last few even-numbered Transformer layers, noting that, despite using fewer MoE layers, the impact on performance is minimal while computational speed is significantly enhanced. DeepSeekMoE-16B/-145B (Dai et al., 2024) replaces all FFNs with MoE layers, excluding the first, due to the observed slower convergence of load balance status in the first layer. MoE-LLaVA (Lin et al., 2024), a recently popular open Large Vision-Language Model (LVLM), demonstrates that alternating MoE placement yields superior model quality and execution efficiency on multimodal tasks, compared to every-layer MoE placement or ”First-Half” and ”Second-Half” configurations. ST-MoE (Zoph et al., 2022) found that adding a dense FFN adjacent to each MoE layer can improve model quality. Brainformers (Zhou et al., 2023) introduce a nonuniform architecture that integrates MoE layers, dense FFNs, attention mechanisms, and a variety of layer normalizations and activation functions without strict sequential layering, trading architectural regularity for the flexibility of sub-layer composition. Jamba (Lieber et al., 2024) integrates the architecture of Mamba (Gu and Dao, 2023) by adopting a 1:7 ratio of attention-to-Mamba layers.
| Name | Time | Affiliation | Params. | Benchmarks | Link | ||||
| Activ. | Total | MMLU | GSM8K | MATH | HumanEval | ||||
| Mixtral-8x7B-v0.1 | 2023.12 | Mistral | 13B | 47B | 70.6 | 58.4, 74.4 (8-shot) | 28.4 | 40.2 | https://huggingface.co/mistralai/Mixtral-8x7B-v0.1 |
| DeepSeekMoE-16B-Base | 2024.1 | DeepSeek | 3B | 16B | 45.0 | 18.8 (8-shot) | 4.3 | 26.8 | https://huggingface.co/deepseek-ai/deepseek-moe-16b-base |
| Grok-1 | 2024.3 | xAI | 86B | 314B | 73.0 | 62.9 | 23.9 | 63.2 | https://github.com/xai-org/grok-1 |
| Qwen1.5-MoE-A2.7B | 2024.3 | Alibaba | 3B | 14B | 62.5 | 61.5 (8-shot) | - | 34.2 | https://huggingface.co/Qwen/Qwen1.5-MoE-A2.7B |
| DBRX Instruct | 2024.3 | Databricks | 36B | 132B | 73.7 | 72.8 | - | 70.1 | https://huggingface.co/databricks/dbrx-instruct |
| Jamba-v0.1 | 2024.3 | AI21 Labs | 12B | 52B | 67.4 | 59.9 (3-shot) | - | 29.3 | https://huggingface.co/ai21labs/Jamba-v0.1 |
| Mistral-8x22B-v0.1 | 2024.4 | Mistral | 39B | 141B | 77.8 | 78.6, 88.4 (8-shot) | 41.8 | 45.1 | https://huggingface.co/mistralai/Mixtral-8x22B-v0.1 |
| Arctic Instruct | 2024.4 | Snowflake | 17B | 480B | 67.3 | 74.2 | - | - | https://huggingface.co/Snowflake/snowflake-arctic-instruct |
| DeepSeek-V2 | 2024.5 | DeepSeek | 21B | 236B | 78.5 | 79.2 (8-shot) | 43.6 | 48.8 | https://huggingface.co/deepseek-ai/DeepSeek-V2 |
| DeepSeek-V2-Chat (RL) | 2024.5 | DeepSeek | 21B | 236B | 77.8 | 92.2 (8-shot) | 53.9 | 81.1 | https://huggingface.co/deepseek-ai/DeepSeek-V2-Chat |
| Yuan 2.0-M32 | 2024.5 | IEIT | 4B | 40B | 72.2 | 92.7 (8-shot) | 55.9 (8-shot) | 74.4 | https://huggingface.co/IEITYuan/Yuan2-M32 |
| Skywork-MoE-Base | 2024.6 | Kunlun | 22B | 146B | 77.4 | 76.1 | 31.9 | 43.9 | https://huggingface.co/Skywork/Skywork-MoE-Base |
4.2.3. Activation Function
Building upon dense Transformer architectures, sparse MoE models have adopted a progression of activation functions paralleling those in leading dense large language models, including BERT (Devlin et al., 2018), T5 (Raffel et al., 2020), GPT (Brown et al., 2020), LLAMA (Touvron et al., 2023) and so on. The evolution of activation functions has seen a shift from ReLU (Glorot et al., 2011) to more advanced options such as GeLU (Hendrycks and Gimpel, 2016), GeGLU (Shazeer, 2020), and SwiGLU (Shazeer, 2020). This trend extends to other components of MoE models, which now frequently incorporate Root Mean Square Layer Normalization (RMSNorm) (Zhang and Sennrich, 2019), Grouped Query Attention (GQA) (Ainslie et al., 2023), and Rotary Position Embeddings (RoPE) (Su et al., 2024).
4.2.4. Shared Expert
DeepSpeed-MoE (Rajbhandari et al., 2022) innovatively introduces the Residual-MoE architecture, wherein each token is processed by a fixed expert and another selected through gating, achieving two experts engagement per layer without increasing the communication cost beyond that of top-1 gating. This approach considers the gating-selected MoE expert as an error-correcting adjunct to the fixed dense FFN. A conceptually similar approach, Conditional MoE Routing (CMR), is employed in NLLB (Costa-jussà et al., 2022), which also combines the outputs of dense FFN and MoE layers. This paradigm of integrating fixed FFN with sparse MoE, often referred to as shared expert and illustrated in Figure 5 (b), has gained traction in recent language models such as DeepSeekMoE (Dai et al., 2024), OpenMoE (Xue et al., 2024), Qwen1.5-MoE (Team, 2024b), and MoCLE (Gou et al., 2023), indicating its ascension to a mainstream configuration. Instead of using a single shared expert, DeepSeekMoE (Dai et al., 2024) and Qwen1.5-MoE (Team, 2024b) employ multiple shared experts, due to their fine-grained expert segmentation design.
However, shared expert configuration, while effective in NLP tasks, have not demonstrated the same level of enhancement in vision tasks. Empirical evidence from ScMoE (Cai et al., 2024) indicates that pairing one shared expert with one gating-selected expert yields only comparable performance to standard top-1 MoE in SwinV2-MoE models. Additionally, based on the design of shared expert, ScMoE decouples the MoE process to separately handle the representations from preceding layers and integrate them with the outputs processed by the shared expert of the current layer, thus improving efficiency by facilitating overlap in communication and computation. A comparable method to enhance overlapping is employed in the Dense-MoE hybrid transformer architecture, as delineated in Snowflake Arctic (Team, 2024c), which bears resemblance to the Lora MoE framework discussed in Section 4.3.3 and illustrated in Figure 6 (d).
4.3. Mixture of Parameter-Efficient Experts
LLMs pretrained on generic massive datasets have demonstrated impressive abilities, enabling their deployment across diverse tasks (Gao et al., 2024). However, to tailor a pretrained LLM for a specific downstream task, fine-tuning is essential. Traditional full fine-tuning, which updates all the parameters of the base model, is computationally intensive, especially as model sizes continue to grow (Ding et al., 2022). To address this issue, research into parameter-efficient fine-tuning (PEFT) has emerged, intending to reduce computational demands during the adaptation of a generic pre-trained model to particular tasks (Han et al., 2024). PEFT methods only update a small set of parameters while maintaining the rest of the base model untouched (Lialin et al., 2023). These techniques have achieved state-of-the-art performance across numerous NLP tasks (Liu et al., 2022; Hu et al., 2021).
Despite these successes, PEFT approaches often struggle with generalizing across multiple tasks due to their limited scope of trainable parameters and the potential for catastrophic forgetting (Li et al., 2024b). To mitigate these limitations, a line of mixture of parameter-efficient experts (MoPE) research has emerged, focusing on integrating the MoE framework with PEFT (Li et al., 2024b; Liu et al., 2023). MoPE incorporates the MoE’s gating mechanism and multi-expert architecture, but with each expert constructed using PEFT techniques (Ostapenko et al., 2023). The subtle combination boosts PEFT’s performance under the multi-task scenario (Zhu et al., 2023b). Additionally, by leveraging PEFT for constructing experts, MoPE operates with fewer parameters, achieving greater resource efficiency compared to traditional MoE models (Zadouri et al., 2023).
MoPE harnesses the best of both fields: the task versatility of MoE and the resource efficiency of PEFT (Li et al., 2024b), positioning it as a promising area of study that pushes the boundaries of both fields. In the following subsection, we will give a taxonomy of MoPE, as depicted in Figure 6, based on their placement within the Transformer model architecture. We will then review recent MoPE research, summarizing the methodologies and contributions of each study.
4.3.1. Feed-Forward Network
Following the conventional MoE structure, a series of studies introduce the MoPE framework to the FFN layer of every Transformer block. During the training process, the focus is on optimizing the parameter-efficient experts and the gating mechanism, leaving the rest of the pre-trained model intact. As illustrated in Figure 6(b), the forward process under the MoPE framework integrated with FFN can be expressed as:
where and is the parameter-efficient expert and gating function applied to the FFN layer, respectively.
One of the pioneering works in this domain, LoRAMoE (Dou et al., 2023b), efficiently applies the MoPE structure to FFN. LoRAMoE integrates a few plug-in LoRA experts into the FFN layer, employing a gating mechanism to orchestrate the experts’ contributions. Realizing the diversity in data distributions, LoRAMoE separates the experts into two distinct groups: one focuses on learning various downstream tasks, and the other is dedicated to aligning pretrained world knowledge with human instructions. To ensure that each group of experts maintains its focus, LoRAMoE defines a localized balancing constraint loss, which preserves the importance of each expert within its group while allowing different groups to concentrate on their respective tasks. This design enables LoRAMoE to effectively resolve the knowledge forgetting issue and enhance model performance on downstream tasks. In a similar vein, AdaMix (Wang et al., 2022) injects a set of Adapter (Houlsby et al., 2019) experts after the FFN layer in each Transformer block. Adapter tuning is a PEFT method that integrates a pair of feed-forward up and down projection matrices into the Transformer block. During fine-tuning, only the incremental Adapter blocks are updated, with the rest of the model unchanged. AdaMix utilizes a stochastic routing policy that randomly selects the projection matrices during training, maintaining computational costs equivalent to a single adapter. To minimize service costs during inference, AdaMix averages the outputs of all experts.
Taking a different approach, MixDA(Diao et al., 2023) includes two training stages to leverage domain-specific knowledge while preserving learned information. During the first stage, MixDA only fine-tunes the domain-adapters that work parallel to the FFN to acquire domain-specific knowledge and keep the world knowledge simultaneously. In the second stage, MixDA introduces a gating network and task-adapters on top of the FFN layer for tailoring the model to specific downstream tasks. This strategy allows for a more nuanced adaptation to the task at hand. LLaVA-MoLE(Chen et al., 2024) extends the application of MoPE to multimodal tasks. It creates a set of LoRA experts for the FFN layer to handle inputs from different domains, enhancing the model’s versatility. LLaVA-MoLE adopts a top-1 routing strategy, activating the most relevant expert based on the router’s output distribution, thus maintaining computational costs close to a standard FFN with LoRA. This framework is effective in addressing data conflicts and consistently surpasses plain-LoRA baselines across diverse data configurations.
Contrasting with the MoPE implementations we have discussed, MixLoRA(Li et al., 2024b) creates a LoRA-MoE framework that closely aligns with the conventional MoE models. Rather than just plugging in multiple lightweight experts, MixLoRA fuses LoRA experts with the shared FFN layer. By leveraging the base weights from a single FFN of the base model, MixLoRA streamlines the creation of the MoPE architecture. Furthermore, MixLoRA implements a high-throughput framework that significantly reduces token computation latency and memory usage during both training and inference, optimizing performance and efficiency.
4.3.2. Attention
A branch of research has been exploring the application of the MoPE framework with the attention mechanism. These studies typically involve augmenting the attention mechanism by incorporating a gating network and a set of parallel experts. The MoPE framework can be applied to the Query, Key, Value, and Output projection modules, individually or in various combinations, within the attention mechanism. During the fine-tuning process, only the parameters of the activated experts and the gating network are updated, while the remaining parameters of the model are kept frozen. For example, as shown in Figure 6(a), the integration of MoPE with the Key and Value module of the attention mechanism can be formalized as follows:
where represents the Query, Key and Value modules, respectively. and denote the parameter-efficient expert and its corresponding gating function for the Key module. Similarly, and indicate the expert and the gating function for the Value module. Here, is the number of experts, and is the dimensions in the Multi-head Attention mechanism.
Recent studies have demonstrated the effectiveness of extending MoE to the attention layer (Zhang et al., 2022b; Shen et al., 2024, 2023c). Additionally, there is a new line of research has focused on the fusion of MoPE with the attention mechanism to enhance the model’s efficiency and adaptability. For instance, MoELoRA (Luo et al., 2024) applies MoE to the attention mechanism in a resource-efficient manner by leveraging a PEFT method. MoELoRA adopts LoRA (Hu et al., 2021) to construct the experts. LoRA introduces two low-rank matrices to receive incremental updates associated with the task-specific fine-tuning. Only the LoRA matrices are updated while the base model is kept untouched during fine-tuning. Specifically, MoELoRA sets multiple LoRA experts to the Query and Value matrices of the attention mechanism, and utilizes a gating network to activate the top experts related to the specific tasks during both training and inference phases. To alleviate routing randomness, MoELoRA employs a contrastive learning loss to control the training of experts. The contrastive learning loss is designed to accentuate the differences in output distributions between experts, thereby encouraging them to capture diverse features relevant to the downstream tasks. MoELoRA offers a solution for flexibly combining various computational modules tailored to downstream tasks.
Another framework, MoCLE(Gou et al., 2023), aims to resolve task conflicts that arise from the diversity of training tasks of different sources and formats. MoCLE utilizes a clustering model to categorize different tasks and then leverages a router to direct the clustered input to LoRA experts inserted into the Query and Value modules of the attention mechanism. These LoRA experts contain a group of multiple task experts and a universal expert. Each task expert is dedicated to a particular task to reduce task conflicts, while the universal expert, trained on all tasks, helps to maintain model generalization. SiRA(Zhu et al., 2023b) introduces several lightweight LoRA adapters as experts, along with a top gating mechanism. To mitigate load imbalance and over-fitting issues, SiRA incorporates a capacity constraint that limits the number of tokens each expert can process. Additionally, it employs an auxiliary loss to promote load balancing and an expert dropout mechanism to equalize the gating distribution. SiRA provides an efficient and fine-grained approach to improving the quality of LoRA.
4.3.3. Transformer Block
The integration of MoPE with the Transformer architecture has received substantial attention in recent research. This approach involves creating two groups of experts: one for the attention mechanism, and another for the FFN within the Transformer block. Each group is regulated by its gating mechanism to control the activation of the experts. As exhibited in Figure 6(c), the forward process under the MoPE framework integrated with the Transformer block can be denoted as:
MoV (Zadouri et al., 2023) is one of the notable attempts that combine MoPE with the Transformer block to pursue parameter efficiency. Utilizing the PEFT method, (Liu et al., 2022), MoV introduces tunable vectors that re-scale the Key and Value modules in the attention mechanism, as well as the activation within the FFN. By substituting conventional experts with vectors and updating only these lightweight experts and their corresponding gating during fine-tuning, MoV significantly reduces the computational burden associated with gradient calculations and lessens the memory footprint required for model storage. Similarly, MoLORA (Zadouri et al., 2023) employs multiple LoRA experts to the attention and FFN blocks, outperforming the standard LoRA approach. UniPELT (Mao et al., 2022) proposed a hybrid framework that integrates three representative PEFT methods as experts, namely Adapter (Houlsby et al., 2019), Prefix-tuning (Li and Liang, 2021), and LoRA (Hu et al., 2021). Prefix-tuning is a method that freezes the base model and optimizes the continuous task-specific vectors prepended to the input of the attention. Within the UniPELT framework, LoRA matrices are applied to the weight matrices of Query and Key in the attention mechanism, Prefix vectors are added to the Key and Value modules, and the Adapter block is inserted after the FFN layer. UniPELT leverages different gating mechanisms to dynamically activate the experts, efficiently finding the approaches that best suit the given task.
Further broadening the scope of the LoRA-MoE framework, Omni-SMoLA(Wu et al., 2023a) extends the MoPE with three sets of LoRA experts, each tailored to handle text tokens, visual tokens, and multimodal tokens, respectively. The specialization enables the architecture to enhance performance across various vision-and-language tasks. In the context of MoPE research, the number of experts emerges as a critical hyperparameter influencing downstream task performance (Zadouri et al., 2023; Wu et al., 2024a). Additionally, the use of many experts will lead to redundancy (Chen et al., 2023b). MoLA (Gao et al., 2024) is one of the pioneering work that explores the expert allocation issue. It proposes a LoRA-MoE framework with a Layer-wise Expert Allocation, which enables the flexible employment of varying numbers of experts across different layers. The expert allocation strategy proposed by MoLA further improves the effectiveness of the LoRA-MoE framework. In the specialized field of medical applications, MOELoRA(Liu et al., 2023) tackles the challenges of task variety and high adaptation cost. It integrates LoRA experts and task-motivated gate functions into the attention and FFN of each layer. The gating utilizes task identity to modulate expert contributions, creating unique parameter sets tailored to individual tasks. The design of MOELoRA combines the strengths of both MOE and LoRA, strengthening LLM’s capability in medical domains.
4.3.4. Every Layer
There has been considerable interest in incorporating MoPE into fundamental components such as the attention, FFN, and Transformer block. Existing work often approaches the attention mechanism and FFN independently, employing distinct gating mechanisms to modulate them separately. Rather than treating these elements isolated, there is a new direction that considers the Transformer layer as an integrated whole. This shift in perspective allows for the application of the MoPE framework to the entire Transformer layer, capturing the combined dynamics of the attention and FFN within a unified approach. As illustrated in Figure 6(d), the forward process under the MoPE framework integrated with every layer is as follows:
where and is the parameter-efficient expert and gating function applied to the entire layer, respectively.
In this context, the approach presented by MoLE(Wu et al., 2024a) provides innovative insights. MoLE identifies that various layers within LoRA exhibit unique features. In response to this finding, MoLE pursues to enhance the composition effect of trained LoRAs by dynamically adjusting the layer-specific weights according to the desired objective. This is achieved by plugging a set of trained LoRAs alongside a gating function into each layer. MoLE treats each layer of trained LoRAs as an individual expert and only trains the gating to learn the optimal composition weights for a specified domain. This dynamic linear composition strategy significantly extends the versatility of LoRA, enabling its application across a broader spectrum of practical scenarios.
4.4. Training & Inference Scheme
The architectural advancements of Mixture-of-Experts (MoE) models have been complemented by extensive research into training and inference schemes, with the objective of optimizing both computational efficiency and model quality.
Original Training & Inference Scheme. Initial training methodologies, as established in seminal works (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022b; Zoph et al., 2022; Riquelme et al., 2021), involve constructing an MoE model and training it from scratch, with inference directly following the model configurations of training.
The advent of MoE models has introduced novel paradigms for training and inference, enabling a flexible approach that synergizes the strengths of dense and sparse models while mitigating their respective weaknesses. As depicted in Figure 7, we divide the emerging schemes into three distinct categories: Dense-to-Sparse, which entails initiating with dense model training and progressively transitioning to a sparse MoE configuration (Wu et al., 2022; Dua et al., 2022; Komatsuzaki et al., 2022; Tianwen Wei, 2024; Lin et al., 2024; Pan et al., 2024; Nie et al., 2021; Chen et al., 2022d); Sparse-to-Dense, which involves degrading a sparse MoE model to a dense form that is more conducive to hardware implementation for inference (Xue et al., 2022a; Chen et al., 2022b; Huang et al., 2023); and Expert Models Merging, a process of integrating multiple pre-trained dense expert models into a unified MoE model (Li et al., 2022; Sukhbaatar et al., 2024; Wang et al., 2023).
4.4.1. Dense-to-Sparse.
To mitigate the training overhead associated with vision MoE transformer models, the Residual Mixture-of-Experts (RMoE) approach (Wu et al., 2022) commences with training a dense, non-MoE model on the upstream task, followed by an efficient fine-tuning stage to transition into a MoE model. This process reveals that directly inheriting expert weights from a pre-trained non-MoE model’s FFN can lead to suboptimal performance, necessitating an alignment between the MoE and FFN outputs during the fine-tuning phase. Similarly, Dua et al. (Dua et al., 2022) advocate for initially training a dense model, subsequently introducing sparsity by incorporating a randomly initialized gating module, and further training the model’s feed-forward layers under sparsity conditions—specifically, by updating the weights locally within each compute node rather than averaging gradients across nodes.
Nie et al. (Nie et al., 2021) present EvoMoE, an efficient end-to-end MoE training framework. EvoMoE decouples the joint learning of experts and the sparse gate, emphasizing the acquisition of foundational knowledge through a single expert at the inception of training. Subsequently, it spawns multiple diverse experts and advances the diversification of experts by training with the novel Dense-to-Sparse gate (DTS-Gate). The DTS-Gate initially operates as a dense activation of all experts, then progressively and adaptively constricting to route tokens to a reduced number of experts. A similar strategy is employed in the development of the MoE-LLaVA (Lin et al., 2024) large vision-language model, which commences with a dense model, subsequently multiplies the feedforward network (FFN) to create expert initializations, and proceeds to train exclusively the MoE layers, while keeping the remaining model components static.
Komatsuzaki et al. (Komatsuzaki et al., 2022) highlight the efficiency of sparse models in terms of quality and computational cost, yet acknowledge their significant data requirements and the expense of training from scratch at scale. To address this, they introduce a scheme termed ”sparse upcycling,” which leverages pre-existing training investments by initializing a sparsely activated MoE model from a pre-trained dense checkpoint. This involves transferring all parameters—and optionally their associated optimizer states—from the original checkpoint, with the exception of the MoE gating network parameters, which are not present in the dense model. Notably, the new MoE layers are populated with identical copies of the original dense model’s FFN layers, and the gating mechanism weights are initialized randomly. A critical obstacle in model upcycling is the initial performance decrease resulting from structural modifications to a trained network. To mitigate this performance regression during upcycling, the researchers proposes normalizing the gate combine weights for each token to 1. This approach is grounded in the notion that, in the dense model, each token was processed by a singular ”expert” FFN. While this normalization proved beneficial for upcycled vision models, it was found to be detrimental to the performance of upcycled language models.
Building upon the sparse upcycling technique (Komatsuzaki et al., 2022), the Skywork-MoE model (Tianwen Wei, 2024) leverages the foundational architecture of its pre-developed Skywork-13B model (Wei et al., 2023), adopting its dense checkpoints as a foundation for initial states. Their empirical evidence indicates that the decision between sparse upcycling and training from scratch should be informed by both the performance of available dense checkpoints and the MoE-specific training resources, as models trained from scratch consistently surpass their upcycled counterparts in performance. The study observes a decline in average expert similarity throughout the training of upcycled MoEs, suggesting a diversification of experts emerges during the process. Importantly, the Skywork-MoE analysis reveals that models with greater expert similarity tend to underperform, establishing expert similarity as a potential diagnostic tool during MoE training for upcycled models. Conversely, the expert similarity in models trained from scratch remains minimal, implying that non-uniform expert initialization promotes diversification.
Pan et al. (Pan et al., 2024) posit that the parameter inefficiency observed in MoE models stems from conventional sparse training methodologies, where only a select group of experts is engaged and refined for each input token. To counteract this, they introduce a hybrid framework for MoE models, denoted as DS-MoE, which integrates dense training (activating all experts) with sparse inference (sparse expert activation) to achieve higher computation and parameter efficiency. Notably, DS-MoE maintains activation for all self-attention experts (MoA (Zhang et al., 2022b)) during inference but selectively activates FFN experts, reflecting the observation that self-attention layers manifest considerably less sparsity compared to FFN layers.
Chen et al. (Chen et al., 2022d) introduce SMoE-Dropout, an innovative plug-and-play training framework, which initially modularizes the FFN into a sequence of smaller FFNs then employs a random policy parameterized by fixed weights to route token to k experts with the largest response. Progressively, the framework activates an increasing number of experts, preventing overfitting to the amounts of used network capacity during training. MoEfication (Zhang et al., 2022a) investigates various strategies for expert construction in T5 models, including random splitting, parameter clustering, and building co-activation graphs. MoEBERT (Zuo et al., 2022) implements an importance-based method for adapting FFNs into experts within BERT models. LLaMA-MoE (Team, 2023) conducts an extensive examination of different expert construction methodologies, ultimately proposing a straightforward random division approach that partitions the parameters of FFNs into non-overlapping experts.
4.4.2. Sparse-to-Dense.
Switch Transformer (Fedus et al., 2022b) studies the distillation of large sparse models into smaller dense counterparts to achieve parameter efficiency for deployment. The study reports that initializing the corresponding weights of dense model from non-expert layers of MoE model modestly enhances performance, facilitated by the consistent dimension of non-expert layers. Furthermore, an improvement in distillation is observed when employing a mixture of 0.25 for the teacher probabilities and 0.75 for the ground truth label. Leveraging both methods, the distillation preserves approximately 30% of the sparse model’s quality gains using only about 5% of the parameters. Similarly, Xue et al. (Xue et al., 2022a) address the challenges of overfitting, deployment difficulty, and hardware constraints associated with sparse MoE models. Drawing inspiration from human learning paradigms, they propose a new concept referred to as ’knowledge integration’ aimed at creating a dense student model (OneS) that encapsulates the expertise of a sparse MoE model. Their framework first implements knowledge gathering, explored through a variety of methods such as summation, averaging, top- Knowledge Gathering, and their Singular Value Decomposition Knowledge Gathering. Then, they refine the dense student model by knowledge distillation to mitigate noise introduced from the knowledge gathering. The OneS model retains 61.7% of the MoE’s benefits on ImageNet and 88.2% on NLP datasets. Further investigations into MoE model distillation are also conducted by other researchers (Rajbhandari et al., 2022; Costa-jussà et al., 2022).
Chen et al. (Chen et al., 2022b) highlight the challenges associated with deploying MoE models on resource-constrained platforms, such as cloud or mobile environments. Observing that only a fraction of experts contribute significantly to MoE fine-tuning and inference, they propose a method for the progressive elimination of non-essential experts. This approach retains the advantages of MoE while simplifying the model into a single-expert dense structure for the target downstream task. Similarly, ModuleFormer (Shen et al., 2023c) applies a comparable pruning technique, removing task-unrelated experts for a lightweight deployment. Huang et al. (Huang et al., 2023) separate the training and inference stages for Vision Transformers (ViTs). They substitute certain FFNs in the ViT with custom-designed, efficient MoEs during training. These MoEs assign tokens to experts using a random uniform partition and incorporate Experts Weights Averaging (EWA) on these MoEs at the end of each iteration. After training, the MoEs are converted back to FFNs through averaging of expert weights, thus reverting the model to its original dense ViT for inference.
4.4.3. Expert Models Merging.
Li et al. (Li et al., 2022) introduce the Branch-Train-Merge (BTM) algorithm, a method for the communication-efficient training of language models (LMs). BTM independently trains a set of expert LMs (ELMs), each tailored to a specific domain within the training corpus, such as scientific or legal text. These ELMs, which operate without shared parameters, can be ensembled or parameter-averaged at inference to coalesce into a singular LM. Expanding on this concept, Sukhbaatar et al. (Sukhbaatar et al., 2024) present Branch-Train-MiX (BTX), designed to combine the strengths of BTM and Mixture-of-Experts while addressing their respective limitations. BTX maintains separate training for multiple expert LLMs, akin to BTM, but subsequently integrates these experts within a unified MoE model. Specifically, it consolidates the FFNs from all ELMs into a singular MoE module at each layer, with a gating network determining the appropriate FFN expert for each token. Other components, such as the self-attention layers from ELMs, are merged by averaging their weights. The resulting model then undergoes MoE-finetuning on all the combined data to enable the gate to effectively mix the FFN experts.
Wang et al. (Wang et al., 2023) point that while the emergence of Foundation Models made it easier to obtain expert models tailored to specific tasks, the heterogeneity of data at test time necessitates more than a single expert. Accordingly, they explore the Fusion of Experts (FoE) challenge, which aims to integrate outputs from expert models that provide diverse but complementary insights into the data distribution, formulating it as an instance of supervised learning.
4.5. Derivatives
Building upon the principles of algorithm design highlighted earlier, numerous studies have drawn inspiration from the Mixture of Experts (MoE) framework, proposing a range of MoE variants. We categorize these innovative models as derivatives of the MoE. For instance, Xue et al. (Xue et al., 2022b) introduced WideNet, an approach that increases model width by substituting the feed-forward network (FFN) with an MoE layer while maintaining shared trainable parameters across Transformer layers, except for the normalization layers. Subsequently, Tan et al. (Tan et al., 2023) presented the Sparse Universal Transformer (SUT), an efficient enhancement of the Universal Transformer, which is characterized by parameter-sharing across its layers. SUT incorporates a Sparse Mixture of Experts and a novel stick-breaking-based dynamic halting mechanism, thus reducing computational complexity without compromising parameter efficiency or generalization capabilities. Moreover, the traditional MoE models often employ discrete matching between experts and tokens (Shazeer et al., 2017; Lepikhin et al., 2020; Du et al., 2022; Artetxe et al., 2021; Zhou et al., 2022a; Fedus et al., 2022b; Zoph et al., 2022), a practice associated with training instability and uneven expert utilization. Addressing these challenges, Antoniak et al. (Antoniak et al., 2023) innovatively proposes the Mixture of Tokens (MoT), which blends tokens from different examples before presenting them to the experts. Thus, MoT enables the model to benefit from a wider array of token-expert combinations.
Recently, the MoE’s principle of assigning specialized knowledge to individual experts has been adapted to parameter-efficient fine-tuning (PEFT). Choi et al. (Choi et al., 2023) propose the sparse mixture-of-prompts (SMoP), a method that utilizes a gating mechanism to train multiple short soft prompts, each adept at processing distinct subsets of data. This addresses the inefficiencies encountered with long soft prompts during prompt tuning. The MoE framework has also been integrated into lifelong learning (LLL), which seeks to facilitate continuous learning from an ongoing stream of data. The Lifelong-MoE model (Chen et al., 2023c) dynamically expands model capacity by adding experts with regularized pretraining, effectively mitigating the issue of catastrophic forgetting (Kirkpatrick et al., 2017) typically associated with straightforward fine-tuning. In a recent development, the MoE concept of conditional computation has been further refined to optimize resource allocation in transformer-based language models (LMs). The Mixture-of-Depths (MoD) (Raposo et al., 2024) employs a binary gating network to decide whether a token should be processed by a given Transformer layer. As a result, MoD transformers can dynamically allocate computational resources (FLOPs) to specific sequence positions, achieving a lower overall FLOP footprint compared to vanilla or MoE-based transformers.
In summary, the evolution of MoE derivatives reveals a trend where models either integrate the conditional computation aspect of the gating mechanism or merge the MoE structure with various tasks achieved by assigning specialized knowledge to individual experts, such as aforementioned prompt tuning (Choi et al., 2023) and lifelong learning (Chen et al., 2023c) with MoE, demonstrating the versatility and adaptability of the MoE architecture across different domains.
5. System Design of Mixture of Experts
While Mixture of Experts (MoE) has been increasingly leveraged to enhance the capabilities of large language models, its adoption introduces new challenges to existing training and inference systems, due to the inherently sparse and dynamic nature of its computational workload. Since the introduction of expert parallelism by GShard(Lepikhin et al., 2020), which implements parallel gating and expert computation by dispatching partitioned local tokens with the load balancing limit of expert capacity, this paradigm has emerged as a fundamental strategy to facilitate the efficient scaling of MoE models. This approach can be viewed as an augmentation of data parallelism (Rajbhandari et al., 2020; Ren et al., 2021; Rajbhandari et al., 2021), where each expert in an MoE layer is assigned to a distinct device, while all non-expert layers are duplicated across devices. As depicted in Figure 8(a), the process flow of expert parallelism consists of the following sequential operations: gate routing, input encode, All-to-All dispatch, expert computation, All-to-All combine, and output decode. In general, the input size for GEMM needs to be large enough to achieve optimal utilization and throughput of computing device necessitates. Therefore, input encode is employed to aggregate the input tokens of a same expert into a contiguous memory space, as determined by the token-expert mapping from gate routing. Subsequently, the All-to-All dispatch is employed to send the input tokens to their corresponding experts across the distributed devices. Following the localized computation by the experts, the inverse process - All-to-All combine and output decode, reinstates the original data layout according to the the gating indices.
Furthermore, the synergy of expert parallelism (Fedus et al., 2022b; Hwang et al., 2023a; Singh et al., 2023; Ma et al., 2022; Zheng et al., 2022) with other existing parallel strategies (tensor (Shoeybi et al., 2019; Smith et al., 2022; Narayanan et al., 2021), pipeline (Huang et al., 2019; Narayanan et al., 2019; Qi et al., 2023), sequence parallelism (Li et al., 2021; Korthikanti et al., 2023; Jacobs et al., 2023)) has been investigated to enhance the scalability and efficiency of MoE models within large-scale distributed environments. As shown in Figure 8, we illustrate several examples pf hybrid parallelism, encompassing (b) data + expert + tensor parallelism (Fedus et al., 2022b; Rajbhandari et al., 2022; Singh et al., 2023; Hwang et al., 2023a; Zhai et al., 2023), (c) data + expert + pipeline parallelism (He et al., 2022; Zhai et al., 2023; Hwang et al., 2023a), (d) expert + tensor parallelism (Tianwen Wei, 2024). It is imperative to recognize that the choice of distributed parallelism strategies influences a complex interplay between computation efficiency, communication overhead, memory occupation, potentially affected by various hardware configurations. Consequently, the deployment strategies for practical applications necessitate nuanced trade-offs and bespoke designs tailored to specific use-case scenarios.
In the subsequent discussion, we delineate the challenges introduced by MoE models from computation, communication, and storage aspects, concurrently reviewing existing research addressing these issues. Table 4 shows an overview of the open-source MoE frameworks.
| Reference | Affiliation | Optimizations | Link | Star | ||
| Computation | Communication | Storage | ||||
| OpenMoE (Xue et al., 2024) | Colossal-AI | ✓ | ✓ | https://github.com/hpcaitech/ColossalAI | 38K | |
| ScatterMoE (Tan et al., 2024) | Mila Quebec | ✓ | https://github.com/shawntan/scattermoe | 140 | ||
| Megablocks (Gale et al., 2023) | Stanford University | ✓ | https://github.com/stanford-futuredata/megablocks | 1.1K | ||
| Tutel (Hwang et al., 2023a) | Microsoft | ✓ | ✓ | https://github.com/microsoft/tutel | 672 | |
| SE-MOE (Shen et al., 2022) | Baidu | ✓ | ✓ | ✓ | https://github.com/PaddlePaddle/Paddle | 21K |
| HetuMoE (Nie et al., 2022) | Peking University | ✓ | ✓ | https://github.com/PKU-DAIR/Hetu | 236 | |
| Deepspeed-MoE (Rajbhandari et al., 2022) | Microsoft | ✓ | ✓ | https://github.com/microsoft/DeepSpeed | 33K | |
| FastMoE (He et al., 2021) | Tsinghua University | ✓ | ✓ | https://github.com/laekov/fastmoe | 1.4K | |
| Fairseq (Ott et al., 2019; Artetxe et al., 2021) | Meta | https://github.com/facebookresearch/fairseq/tree/moe | 29K | |||
| Mesh-TensorFlow (Shazeer et al., 2018) | https://github.com/tensorflow/mesh | 1.6K | ||||
5.1. Computation
Despite MoE is designed to scale model parameters efficiently without increasing computational demand, it encounters challenges pertaining to computational efficiency. One concern is the imbalance of computational load across distributed devices employing expert parallelism, which incurs significant synchronization overhead as the system awaits the processing completion of the most heavily loaded expert. Such issues are typically addressed through algorithmic strategies, such as optimized gating mechanisms and expert capacity adjustments, as discussed in the preceding section. Besides, solutions like SE-MoE (Shen et al., 2022), Tutel (Hwang et al., 2023a), FlexMoE (Nie et al., 2023) and SmartMoE (Zhai et al., 2023) have introduced dynamic expert placement strategies to distribute the workload as equally as possible among devices. Additionally, FasterMoE (He et al., 2022) has implemented a novel dynamic shadowed expert strategy, replicating experts on multiple devices to mitigate severe load imbalance. These model placement related strategies impact both computation and communication efficiency.
Another concern is that MoE introduces additional computational overhead through operations including gate routing, input encode and output decode. Unlike expert computations, which mirror operations in dense models and benefit from extensive optimization on prevalent hardware such as GPUs, these MoE operations are characterized by redundant computation and memory movement, resulting in low efficiency on computing devices. Therefore, recent studies like DeepSpeed-MoE(Rajbhandari et al., 2022), FastMoE (He et al., 2021), HetuMoE (Nie et al., 2022) and Tutel (Hwang et al., 2023a) have focused on the development of tailored GPU kernels to enhance the efficiency of MoE operations.
In contexts where multiple experts are deployed on a single GPU device, MegaBlocks (Gale et al., 2023) reformulates MoE computation in terms of block-sparse operations, developing specialized block-sparse GPU kernels that efficiently handle the dynamic workloads without dropping tokens. Zheng et al. (Zheng et al., 2023) propose PIT, a deep-learning compiler tailored for dynamic sparsity of MoE, which can find feasible PIT rules for all the operators within a model and generate optimized GPU kernels for them. PIT employs a novel tiling mechanism, utilizing the Permutation Invariant Transformation (PIT)—a mathematically proven property-to transform multiple sparsely located micro-tiles into a GPU-efficient dense tile without changing the computation results, thus achieving both high GPU utilization and low coverage waste. Despite these advancements, Tan et al. (Tan et al., 2024) highlight remaining optimization potential within current MoE frameworks such as MegaBlocks and PIT, which commence with an initial scatter-to-group data copy that increases memory footprint and requires a translation of the MoE problem into the sparse matrix format. Although this translation contributes minimally to computation overhead, it imposes limitations on the transparency and adaptability of extending MegaBlocks to modules beyond the FFN. To address these issues, Tan et al. (Tan et al., 2024) propose ScatterMoE, a MoE implementation designed to effectively minimize the memory footprint. ScatterMoE leverages ParallelLinear, a linear module capable of executing grouped matrix operations on scattered groups. This approach yields intermediate representations (e.g., the hidden states of an SMoE MLP) that are directly accessible as standard PyTorch tensors, allowing for easy extensions of MoE methods to other types of expert modules.
5.2. Communication
In expert parallelism, the quadruple invocation of All-to-All communication during both the forward and backward propagation phases within each MoE layer causes a significant overhead, even emerging as the primary constraint on efficiency. The All-to-All communication paradigm encompasses both intra-node (via PCIe, pre-4th-generation NVLink) and inter-node (Ethernet, Infiniband, 4th-generation NVLink) communication channels. The efficiency of such communication is contingent upon a multitude of factors, including the heterogeneity of channel bandwidths, network topology, and the collective communication algorithms. Moreover, load imbalances intrinsic to MoE may exacerbate these inefficiencies by inducing synchronization delays.
To optimize the use of high intra-node bandwidth and low inter-node bandwidth, DeepSpeed-MoE (Rajbhandari et al., 2022) and HetuMoE (Nie et al., 2022) have introduced a hierarchical All-to-All communication strategy that enhances intra-node process and reduces inter-node data exchanges. Besides, FasterMoE (He et al., 2022), TA-MoE (Chen et al., 2022c) and SE-MoE (Shen et al., 2022) have introduced topology-aware routing strategies aimed at mitigating cross-node expert selection, thereby reducing inter-node communication burdens. Additionally, ExFlow (Yao et al., 2024) exploits expert affinity, anticipating expert allocation across layers to maximize the retention of token processing within local GPU confines. The strategic allocation of experts to minimize network traffic and leverage high-bandwidth connections is a prevalent approach in distributed MoE system (Rajbhandari et al., 2022; Singh et al., 2023; Tianwen Wei, 2024). And this is often integrated with the placement design of non-expert modules to optimize overall system performance.
Given the concurrent feature of communication and computation, pipelining (Huang et al., 2019; Narayanan et al., 2019; Qi et al., 2023) is commonly employed to overlap their execution, thereby reducing the total time cost. This technique, which is integrated in systems such as Tutel (Hwang et al., 2023a), FasterMoE (He et al., 2022), and MPipeMoE (Zhang et al., 2024), orchestrates overlapping between All-to-All communication and expert computation. Notably, Lancet (Jiang et al., 2024b) underscores the inherent constraints of these pipelining methods, particularly the bounded duration for which expert computation and communication can overlap. To address this limitation, Lancet partitions non-MoE computations and integrates them into the pipeline during forward pass, and strategically schedules gradient weight computations to augment overlap in the backward pass. With the same objective of extending the overlap duration, ScMoE (Cai et al., 2024) restructures the MoE architecture to simultaneously process representations from preceding layers while engaging with current-layer representations. This decoupling of communication dependencies facilitates substantial, and in certain cases, complete overlapping between communication and computation. Snowflake Arctic (Team, 2024c) employs a similar design, utilizing a Dense-MoE hybrid transformer architecture to effectively overlap communication with computation.
5.3. Storage
The ever-increasing parameters in MoE models exacerbate the constraints posed by memory capacity in compute devices, a challenge already pronounced in dense models. While expert parallelism offers a mitigation strategy through the distribution of experts across multiple devices, individual devices may still struggle to accommodate numerous experts, particularly in inference contexts where device capacity—such as that of edge devices (PCs, smartphones, IoTs)—is inherently more restricted.
Considering the hierarchical storage pyramid, solutions like SE-MoE (Shen et al., 2022), Pre-gated MoE (Hwang et al., 2023b), and EdgeMoE (Yi et al., 2023) selectively retain only essential non-expert parameters and the active expert parameters within the GPU’s High-Bandwidth Memory (HBM), offloading inactive expert parameters to CPU memory or SSDs. These patterns incur additional overhead from data transfer across the storage hierarchy, thus they integrate expert selection forecasting and expert parameter prefetching techniques to overlap parameter access with computation.
In addition, MPipeMoE (Zhang et al., 2024) introduces a strategy to reduce the memory overhead associated with activations and temporary buffers. This is achieved by sharing buffer for various partitions of tensors, while leveraging recomputation/communication and CPU offloading to recover the requisite activations in the backward pass.
6. Applications of Mixture of Experts Models
In the current landscape dominated by Transformer-based large language models (LLMs), the mixture of experts (MoE) paradigm offers a compelling method to significantly expand model capacity while avoiding a corresponding surge in computational demands during training and inference phases. These models have been instrumental in enhancing the performance of LLMs across a spectrum of downstream tasks, with some applications achieving results that eclipse human performance (Fedus et al., 2022a; Jiang et al., 2024a; Dai et al., 2024). Rumors suggest that the formidable GPT-4 may employ an MoE architecture with an array of B experts, trained on diverse datasets and tasks, and utilizing a 16-iteration inference process 111https://x.com/soumithchintala/status/1671267150101721090. Given these, MoE has garnered widespread adoption across fields such as natural language processing, computer vision, recommender systems, and multimodal applications. The essence of these applications lies in leveraging conditional computation to significantly boost the number of model parameters, thereby augmenting model capacities with a fixed computational cost, or implementing dynamic expert selection through gating mechanisms for efficient multi-task learning. In the following, we will explore several representative applications of MoE in various domains, to provide a overall understanding of how MoE can be utilized to specific tasks.
Natural Language Processing. The integration of MoE architectures with LLMs has unlocked extraordinary capabilities in a range of natural language understanding (NLU) and generation (NLG) tasks, including machine translation (Shazeer et al., 2017; Costa-jussà et al., 2022), open-domain question answering (Du et al., 2022; Artetxe et al., 2021), code generation (Jiang et al., 2024a; Dai et al., 2024; Team, 2024a; Tianwen Wei, 2024), and mathematical problem-solving (Dai et al., 2024; Jiang et al., 2024a; DeepSeek-AI, 2024; Team, 2024a). The methods of integrating MoE into LLMs have been thoroughly discussed and analyzed in the preceding algorithm design section 4 and system design section 5, and will not be reiterated in depth here.
Computer Vision. The great success of sparsely-gated Mixture of Experts networks (MoE) in NLP has inspired their application in computer vision. For example, Riquelme et al. (Riquelme et al., 2021) introduced Vision MoE (V-MoE), which incorporates a sparsely activated mixture of MLPs into select ViT (Dosovitskiy et al., 2020) blocks. In image recognition tasks, V-MoE rivals the performance of state-of-the-art networks while requiring substantially less computational power during inference. This demonstrates the potential of MoE to discern distinct image semantics through specialized experts. Hwang et al. (Hwang et al., 2023a) develop Tutel, a scalable stack design and implementation for MoE with dynamic parallelism and pipelining, which they demonstrate with SwinV2-MoE, built upon Swin Transformer V2 (Liu et al., 2021). Moreover, Zhang et al. (Zhang et al., 2023) explore adversarial robustness in CNN-based MoE models, proposing a novel router-expert alternating adversarial training framework called ADVMOE. In most recent work, Chowdhury et al. (Chowdhury et al., 2023) introduce the concept of patch-level routing in MoE (pMoE) that segments each input image into patches (or tokens) and allocates patches () to each expert for processing through prioritized routing to enhance efficiency.
Recommender System. Recommender systems are quintessential in various large-scale applications where they are required to balance and optimize multiple objectives simultaneously (Zheng and Wang, 2022). A prime example is in the domain of movie recommendations, where the aim is not only to suggest movie that align with users’ immediate preferences but also to ensure subsequent user satisfaction for the selected movies (Ma et al., 2018). The effectiveness of multi-task models hinges on the intricate interplay between task-specific goals and the relationships between tasks. Consequently, understanding the trade-offs inherent in these relationships is crucial. Mixture-of-experts (MoE) models with gating mechanisms have emerged as a popular paradigm for tackling the complexities of multi-task learning in recommender systems. Ma et al. (Ma et al., 2018) introduce the multi-gate mixture-of-experts (MMOE) approach, which capitalizes on the concept of shared expert submodels across all tasks, guided by a gating network tailored to each individual task. Addressing the “seesaw phenomenon” where the improvement of one task’s performance can detrimentally affect another is another challenge in multi-task learning. To counteract this, Tang et al. (Tang et al., 2020) propose the Progressive Layered Extraction (PLE) model for personalized recommendations. PLE distinctly segregates shared and task-specific components and employs a progressive routing mechanism to incrementally extract and refine the semantic knowledge, thereby enhancing the efficacy of joint representation learning and the routing of information across tasks. Recently, in the pursuit of capturing both the long-term and short-term user preferences that are particularly salient in sequential recommendation scenarios, a novel method named AdaMCT (Jiang et al., 2023b) has been proposed. AdaMCT utilizes layer-aware adaptive mixture units to dynamically blend CNN and Transformer experts, thereby tailoring the recommendations to individual user patterns.
Multimodal Applications. Multimodal models are designed to process and integrate various data types within a single neural network framework (Ngiam et al., 2011). These models often simultaneously encompass two primary data modalities: images and text (Baltrušaitis et al., 2018; Uppal et al., 2022; Zhou et al., 2020). The Mixture of Experts (MoE) architecture has gained considerable traction as the foundation of multimodal models due to its capacity for expert layers to learn distinct modality partitioning (Mustafa et al., 2022). One notable implementation of this approach is the LIMoE model (Mustafa et al., 2022), a sparse mixture of expert models tailored for multimodal learning. LIMoE is trained on both images and text data, employing contrastive loss and an entropy-based regularization technique to address load balancing challenges inherent in MoE systems. Subsequently, Shen et al. (Shen et al., 2023b) and Lin et al. (Lin et al., 2024) have further investigated the potential of MoE for scaling vision-language models, offering valuable insights that contribute to the development of more efficient and effective multimodal learning systems. Furthermore, to address the specific issue of task conflicts in instruction tuning of Large Vision-Language Models (LVLMs), MoCLE (Gou et al., 2023) integrates MoE with LoRA (Hu et al., 2021) experts and a distinct universal expert to activate task-specific model parameters based on clusters of instructions. In parallel, to mitigate data conflicts, LLaVA-MoLE (Chen et al., 2024) deploys a set of LoRA experts, specifically for the MLP layer, combined with a top-1 gating mechanism to refine instruction tuning in Multimodal Large Language Models (MLLMs). While the MLLMs employing MoE architectures have demonstrated impressive performances, they generally involve a limited number of experts and modalities (Li et al., 2024a). To address this limitation, Li et al. (Li et al., 2024a) introduce the pioneering Uni-MoE, a unified MLLM with MoE architecture capable of managing an extensive range of modalities. They introduce a progressive training strategy to bolster expert collaboration and generalization across modalities, and they utilize LoRA (Hu et al., 2021), a lightweight fine-tuning methodology, to minimize computational demands.
7. Challenges & Opportunities
Mixture of Experts (MoE) models present a compelling approach for significantly increasing model capacity at a constant computational cost. Despite their promise, several intrinsic challenges remain. In this section, we identify critical challenges and promising directions for future investigation as follows:
Training Stability and Load Balancing. MoE models that utilize sparse gating have become a popular means to expand model capacity without proportionally increasing computational demands. However, the discrete nature of assigning a fixed number of experts to tokens leads to significant challenges in maintaining balanced workloads of experts and training stability across varying inputs (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022b; Zhou et al., 2022a; Antoniak et al., 2023). Load imbalances, where certain experts become over-utilized while others are underutilized can hinder expert specialization and further degrade model performance. Although current efforts (Lepikhin et al., 2020; Fedus et al., 2022b; Du et al., 2022; Jiang et al., 2024a; Lieber et al., 2024; Dai et al., 2024; Tianwen Wei, 2024) have attempted to address this challenge by incorporating auxiliary loss functions to encourage even token distribution across experts, these solutions can still lead to training instability (Zoph et al., 2022) and often neglect the relative importance of different tokens (Zhou et al., 2022a). Therefore, future studies should focus on more effective regularization techniques (Zoph et al., 2022) or innovative gating algorithms (Zhou et al., 2022a; Antoniak et al., 2023; Muqeeth et al., 2023; Zhong et al., 2024) that encourage equitable load distribution among experts and enhance model training stability.
Scalability and Communication Overhead. As the escalating sizes of LLMs with MoE necessitate more expansive distributed systems, the imperative for efficient communication during model training becomes increasingly critical, as elaborated in Section 5.2. The trade-off between model complexity, indicated by the number of parameters, and the communication overhead represents a significant bottleneck in distributed training processes (Lepikhin et al., 2020). To address these challenges, it is essential to develop and implement effective strategies that enhance the efficiency of information transfer from system aspect or streamline information exchange without compromising model performance from algorithm aspect. Innovations such as DeepSpeed (Rajbhandari et al., 2022), FasterMoE (He et al., 2022), and ScMoE (Cai et al., 2024) are at the forefront of minimizing communication overhead. For example, the shared expert approach (Rajbhandari et al., 2022; Dai et al., 2024; Team, 2024b, c), advancing MoE with parameter-sharing frameworks, holds promise for reducing the volume of data transmitted between distributed systems while concurrently enhancing model performance in natural language processing tasks. Such innovations are pivotal in facilitating more scalable and efficient distributed training architectures for MoE models.
Expert Specialization and Collaboration. Expert specialization refers to the concept where each expert develops non-overlapping and focused knowledge. Encouraging experts to concentrate their skills on distinct sub-tasks or domains has been shown to enhance the performance and generalization of the MoE model. The prevailing strategy involves designating a select number of experts as shared ones, with the goal of capturing commonalities in knowledge and reducing redundancy among those experts that are routed dynamically (Rajbhandari et al., 2022; Xue et al., 2024; Team, 2024b; Dai et al., 2024). However, fostering effective collaboration among these specialized experts is an ongoing challenge. Relying solely on a sparsely computed weighted sum of outputs from the top- experts can overlook the intricate internal relationships that exist across the entire experts. Consequently, exploring new mechanisms for enhancing both the specialization and collaboration among experts is crucial for the development of more integrated and powerful MoE models.
Sparse Activation and Computational Efficiency. One of the primary benefits of MoE models lies in their capacity for sparse activations, which theoretically enhances computational efficiency. Nevertheless, implementing this efficiency in practice poses substantial challenges. This is attributed to the non-uniformity of sparse operations within hardware accelerators (Dao et al., 2022; Kwon et al., 2023). Furthermore, optimizing the balance between activating a select top- subset of experts from an entire pool of experts entails intricate coordination. This optimization is crucial for ensuring that each expert develops a specialized niche (Dai et al., 2024). Thus, there is a pressing need for further research into hardware optimization techniques that more adeptly accommodate sparse computations. Such advancements would not only preserve the model’s capacity but could also significantly enhance the performance and efficiency of MoE models.
Generalization and Robustness. MoE models have demonstrated increased computational efficiency during pre-training phases. However, there is a notable propensity for sparse MoE architectures to overfit to specific tasks or datasets, which undermines their ability to generalize effectively (Fedus et al., 2022b; Zoph et al., 2022; Shen et al., 2023a; Dou et al., 2023a). To enhance the generalization and robustness of MoE models when encountering unseen data and diverse input variations, various strategies have been explored. These include regularization techniques such as dropout (Fedus et al., 2022b) and token dropping (Zoph et al., 2022), as well as multi-task instruction tuning (Shen et al., 2023a; Dou et al., 2023a). Looking ahead, there is potential for further advancements in this challenge. Future endeavors could explore innovative regularization methods, refined multi-task learning frameworks, or the incorporation of meta-learning concepts that bolster the MoE models’ robustness and extend their generalization capabilities across an even broader spectrum of downstream tasks.
Interpretability and Transparency. The inherent complexity of MoE models, coupled with their dynamic gating of inputs to specialized experts, poses significant challenges to interpretability. This becomes particularly problematic in contexts where comprehending the rationale behind the model’s decisions is essential. Enhancing the interpretability of MoE models is therefore critical, not only to facilitate a clearer understanding of their decision-making processes but also to address underlying challenges such as load balancing (Lepikhin et al., 2020; Fedus et al., 2022b; Du et al., 2022) and the mitigation of knowledge redundancy (Team, 2024b; Dai et al., 2024). In light of these considerations, there is a pressing need for future studies focused on the development of methods and tools that can effectively visualize and explain the behavior of individual experts within MoE models, as well as the nature of their interactions. Such advancements would significantly improve our grasp of MoE models and bolster their ongoing development, ensuring their gating decisions are transparent and trustworthy.
Optimal Expert Architecture. The design of MoE architectures, encompassing the selection of network types and the quantity of experts, significantly influences the efficacy of multi-task learning across various domains. A plethora of network architectures has been adopted as experts, including LSTM (Shazeer et al., 2017), CNN (Chowdhury et al., 2023; Zhang et al., 2023), FFNs (MLPs) (Lepikhin et al., 2020; Fedus et al., 2022b; Zoph et al., 2022; Pan et al., 2024), Attention (Zhang et al., 2022b; Shen et al., 2023c), and LoRA (Dou et al., 2023b; Li et al., 2024b; Luo et al., 2024). Among these, FFNs as experts remain the most prevalent. Despite their considerable achievements, the exploration of various hybrids of network types within experts (as the distinct features processing capabilities of different network architectures), as well as the development of innovative expert architectures, remains nascent areas of research. Furthermore, the strategic allocation of a varying number of experts across different layers of the model presents an area ripe for investigation. This is due to two primary considerations: 1) different layers of the model capture semantic information at varying levels of granularity; 2) an excessive number of experts can complicate the training process and augment computational costs, while an insufficient number of experts might lead to knowledge redundancy and diminish the specialization capabilities of the experts. To navigate these challenges, the development of automated architecture search methods specifically designed for MoE models is imperative (Zhou et al., 2023). Such approaches could systematically identify optimal configurations, balancing the trade-offs between computational efficiency and the specialization of experts.
Integration with Existing Frameworks. Ensuring seamless integration of MoE models into existing large language models (LLMs) is crucial for their broad adoption. It is particularly vital to enable adaptation of LLMs to MoE architecture without necessitating training from scratch, as it can significantly reduce resource consumption. Recent studies (Dou et al., 2023b; Chen et al., 2024; Li et al., 2024b; Luo et al., 2024; Zadouri et al., 2023; Gao et al., 2024; Wu et al., 2024a) have demonstrated the efficacy of combining Parameter-efficient Fine-tuning (PEFT) techniques with MoE frameworks, offering a promising method for incorporating MoE into established LLMs. However, these methods may compromise model performance or complicate the existing parallel strategies of pretraining and inference efforts (Han et al., 2024). Advancing the development of modular and plug-and-play MoE components is essential. Additionally, optimizing these components for training and deployment across diverse computing environments and hardware platforms will expand their applicability. Such advancements are expected to enhance the versatility and efficiency of MoE models, making them more accessible for a wide range of applications and platforms.
By addressing these challenges, we can unlock the full potential of MoE models, paving the way for more efficient and powerful machine learning systems, particular for large language models (LLMs), that are capable of handling the ever-growing complexity and diversity of real-world tasks.
8. Conclusion
In this survey, we present a systematic and comprehensive review of the literature on MoE models, serving as a valuable compendium for researchers exploring the landscape of MoE technologies. We introduce a new taxonomy for MoE models and provide an in-depth analysis that encompasses three distinct vantage points: algorithm design, system design, and practical applications, complemented by a curated collection of open-source implementations, detailed hyperparameter configurations, and thorough empirical assessments. Moreover, we highlight the critical challenges faced in the field and outline the most promising avenues for future investigation. To support the continuous dissemination of knowledge and advancements, we have established a dedicated resource repository to facilitate ongoing updates and the sharing of cutting-edge developments in MoE research. We hope this survey can contribute to an essential reference for researchers seeking to rapidly acquaint themselves with MoE models, and that it will actively contribute to the vibrant progression.